### v2

# leftmargin=1em, rightmargin=1em, indentfirst=false

Abstract
Large language models (LLMs) have emerged as a dominant AI paradigm due to their exceptional text understanding and generation capabilities. However, their tendency to generate inconsistent or erroneous outputs challenges their reliability, especially in high-stakes domains requiring accuracy and trustworthiness. Existing research primarily focuses on detecting and mitigating model misbehavior in general-purpose scenarios, often overlooking the potential of integrating domain-specific knowledge. In this work, we advance misbehavior detection by incorporating domain knowledge. The core idea is to design a general specification language that enables domain experts to customize domain-specific predicates in a lightweight and intuitive manner, supporting later runtime verification of LLM outputs. To achieve this, we design a novel specification language ESL and introduce a runtime verification framework RvLLM to validate LLM output against domain-specific constraints defined in ESL. RvLLM operates in two main stages: interpretation and reasoning. During interpretation, it derives interpretations of the specification based on context, which then guide the reasoning process to identify inconsistencies. When new knowledge is derived, RvLLM issues a follow-up query to the LLM to further verify the consistency. We evaluate RvLLM on three representative tasks: violation detection against Singapore Rapid Transit Systems Act, numerical comparison, and inequality solving. Experimental results demonstrate that RvLLM effectively detects erroneous outputs across various LLMs in a lightweight and flexible manner. The results reveal that despite their impressive capabilities, LLMs remain prone to low-level errors due to limited interpretability and a lack of formal guarantees during inference, and our framework offers a potential long-term solution by leveraging expert domain knowledge to rigorously and efficiently verify LLM outputs.
1 Introduction
Unlike rule-based systems [47] operating on predefined and deterministic rules, large language models (LLMs) [3, 21, 56, 39] learn data representation and processing automatically from the training datasets, achieving human-like or even superhuman performance, and have driven significant advancements in various practical applications [11, 30, 50, 40, 22]. However, their non-deterministic and unpredictable nature sometimes leads to inconsistent or erroneous outputs [5, 45], posing significant risks in safety-critical or knowledge-intensive domains. Recent studies [55, 6] have shown that such misbehavior in LLMs cannot be fully eliminated, underscoring the urgent need for dedicated verification and validation methodologies to enhance the reliability of LLM-generated outputs.
LLM testing [59, 25, 28, 60] primarily aims to establish comprehensive benchmarks to evaluate the overall model performance against domain-agnostic criteria–such as accuracy, coherence, and fairness–in alignment with the intended application. While these approaches effectively assess general behavior and reveal edge cases that may provoke unexpected responses, they are limited to predefined benchmarks and lack the specificity needed to address domain-specific assessment needs. LLM verification, instead, may serve as a complementary mechanism to LLM testing by providing formal guarantees on model behavior. Despite substantial progress over the past decade in neural network verification [9, 17, 27, 31, 54, 44, 57], existing methods exhibit notable limitations: they are primarily tailored to simpler model architectures–such as deep or convolutional networks–and classification tasks, struggling to scale to the complexity of modern AI models and their diverse functionalities. Furthermore, these approaches target at general properties such as robustness and fairness, making them ill-suited for domain-specific properties. Therefore, developing a dedicated verification paradigm tailored to LLMs is crucial for ensuring certified and reliable outputs.
On the other hand, the decline of rule-based expert systems [23, 29, 20] in the late 20th century can be attributed to their inability to handle incomplete information and poor scalability in real-world complexity. Defining comprehensive rule sets for open domains has been proven infeasible [47, 26]. We argue that such rules can only be statistically approximated–a capability exemplified by LLMs. However, this approximation cannot substitute for explicit rule encoding, and learning-based models remain inherently unreliable, particularly in edge cases demanding strict adherence to domain-specific requirements. This highlights the need for verifying explicitly defined rules in a lightweight and incremental manner for practical deployment and maintainability.
Figure 1: An overview of RvLLM. Given the LLM’s context and outputs, RvLLM first extracts relevant propositions and translates the rules from the specification file into a normalized set of propositional logic formulae. These formulae are then processed via forward chaining to detect inconsistencies or infer new knowledge, which is then used to generate follow-up queries and assess consistency across successive LLM responses.
This paper. Building on these foundations, we propose runtime verification with domain knowledge as a sustainable solution to ensuring reliable LLM behavior. Our work is motivated by two key insights: i) Runtime verification offers a lightweight yet rigorous means to bridge testing and formal verification by assessing system behavior against formal properties during execution; ii) Existing work primarily focuses on general misbehavior detection, and often overlooks the critical role of domain-specific expertise. Integrating such knowledge is essential for handling edge cases and improving LLM reliability in specialized tasks, as domain knowledge can significantly enhance model performance.
To achieve this, we introduce RvLLM, an innovative runtime verification framework tailored to LLMs to ensure reliable outputs, particularly against axiomatic domain specifications. The idea is to automatically check whether the LLM-generated responses adhere to domain-specific properties expressed in a simple, adaptable specification language. This simplicity and flexibility empower domain experts to encode their knowledge and formally define constraints that capture the expected behavior of LLMs in specialized applications. To support this, we introduce ESL, a general language tailored for encoding rule-based domain expertise. ESL integrates natural language with formal logic to ensure the adaptability to the natural language settings of LLMs while enabling the structured and formal representation of properties to verify. This design allows for the rigorous specification of domain-specific criteria across diverse LLM applications.
Figure 1 presents an overview of our proposed framework. Given an ESL specification provided by domain experts, RvLLM first extracts relevant information from the LLM’s context and outputs to interpret the specification, generating a set of propositional formulae along with truth assignments for the corresponding propositions. Following a normalization step, these formulae are transformed into a standard form and subsequently validated using a forward chaining procedure. This process either detects inconsistencies due to logical contradictions or infers new knowledge. Once new knowledge is inferred, the query generation module issues targeted queries, evaluates the LLM responses, and detects inconsistencies when outputs contradict the previously inferred knowledge.
Experimental results. We evaluate the effectiveness of RvLLM through experiments on three representative tasks: violation detection against Singapore Rapid Transit Systems Act [1], numerical comparison, and inequality solving, using a diverse set of LLMs. The experimental results show that RvLLM significantly enhances the reliability of LLM output in these domain-specific tasks. In the violation detection task, employing RvLLM as a complementary mechanism increases the true positive rate, or TPR, by 15.7% to 50.2% across various models. In the numerical comparison task, RvLLM detects nearly all errors in the LLM responses across 100 randomly generated questions. In the inequality solving task, RvLLM facilitates a systematic analysis of the step-by-step reasoning process from LLMs, effectively identifying its deviations from expert-defined specifications. These specifications are primarily grounded in fundamental algebraic inequality properties typically covered at the junior college level. Notably, despite the incomplete domain knowledge in the inequality solving task, RvLLM still achieves a TPR of 52.2% in detecting erroneous solutions. We argue that such a runtime verification approach can function as a critical safeguard for LLMs, particularly in rule-based or reasoning-intensive domains where adherence to specific constraints is essential.
2 Related work
LLM testing. Research in this area aims to develop comprehensive benchmarks and systematic evaluation frameworks to assess LLM across general attributes such as robustness, fairness, and factual consistency [41, 37, 52]. Robustness studies examine model resilience to adversarial input, paraphrasing, or noise [10, 51, 58], while fairness evaluations assess bias with respect to sensitive attributes like gender or race [35, 58]. Factual consistency detection, or hallucination detection, focuses on identifying instances in which LLMs produce plausible yet factually incorrect or unsupported claims [38, 34, 36]. Although large-scale benchmarks like HELM [37] and BIG-bench [49] offer task suites and metrics for rigorous performance evaluation, spanning accuracy, calibration, and ethical considerations, a critical challenge persists, i.e., the dynamic and open-ended nature of real-world deployments complicates the creation of exhaustive and representative test suites.
Runtime verification of LLMs. Runtime verification [7, 19] is a dynamic analysis technique that ensures certified system behavior during execution by monitoring traces against formal specifications. Owing to its lightweight nature, runtime verification has inspired multiple efforts to verify LLM-generated outputs at runtime [14, 8, 43, 12, 33, 42, 32]. However, existing approaches center on open-domain settings, where specifications are often ambiguous and informal. For instance, SelfCheckGPT [43] and CHECKEMBED [8] operationalize specifications in terms of output stability. More recently, [13] proposes a fairness monitoring approach leveraging precisely defined properties in formal specifications. These specifications, informed by linear temporal logic [46] and its bounded metric variant [4], signal a shift towards formal methods for more accurate and dependable runtime verification of LLM behaviors. However, these works primarily focus on general properties and are not well suited to domain-specific constraints for specialized tasks.
3 ESL: a simple way of specifying domain-specific properties
In this section, we introduce a general language, Expert Specification Language (ESL), which can be customized with domain-specified predicates by experts to impose behavioral constraints on LLMs.
3.1 Design guidance
In the following, we give our high-level requirements for designing a specification language that encapsulates LLM behavioral constraints from the perspective of domain experts. The core objective is to strike a balance between user-friendliness and expressiveness. We illustrate this with the following regulation from Singapore Rapid Transit System Act [1]: {quoting} “No person shall consume or attempt to consume any chewing gum or bubble gum while in or upon any part of the railway premises.”
Accessibility for domain experts. The primary goal of this specification language is to empower domain experts, even those with limited programming or formal methods experience, to effectively define and enforce constraints on the LLM’s behavior. To balance expressiveness and user-friendliness, we base the specification rule on predicate logic [48, 18], restricted to prenex normal forms [48, 18] and consisting solely of universal quantifiers. We further constrain its quantifier-free part to a rule-like form (or implication), yielding a simplified variant of predicate logic expressible as $\forall\vec{x}.(f(\vec{x})\Rightarrow g(\vec{x}))$ [^1] We use $\vec{x}$ instead of $x$ as there could be a sequence of variables.. Specifically, both the left-hand side (LHS) and right-hand side (RHS) of the rule-like part are quantifier-free predicate formulas composed of predicates applied to terms that may include variables (e.g., $x$), constants (e.g., ‘bubble gum’), and functions of terms (e.g., $\text{sin}(x)$). For example, ${\sf ChewGum}(x)$ is an atomic predicate with a single variable argument $x$, and ${\sf IsGreater}({\sf Square}(x),0)$ is also an atomic predicate with two arguments: the function term ${\sf Square}(x)$ and the constant term 0.
Ease of customization. The specification language should be designed to support seamless customization, enabling adaptation to diverse domain-specific requirements. To facilitate this, we require each predicate in the specification to be associated with a natural language description that aligns with domain expertise. These descriptions allow customized predicates to remain clear, semantically grounded, and adaptable to specific application contexts. Additionally, built-in terms such as arithmetic functions can be provided to streamline constraint formulation and reduce the need for extensive customization. For example, a domain expert can define their specialized atomic predicates ${\sf ChewGum}(x)$ and ${\sf InRailway}(x)$ associated with a natural language description as follows:
{quoting}
[…]
Efficient interpretation. The language should also support an efficient interpretation procedure to obtain propositional formulae from the specification rule for subsequent logical verification. To this end, we remove the universal quantifier from the standard predicate formulas. For example, the regulation previously mentioned, typically expressed as $\forall x.({\sf InRailway}(x)\Rightarrow\neg{\sf ChewGum}(x))$, will be simplified to ${\sf InRailway}(x)\Rightarrow\neg{\sf ChewGum}(x)$.
3.2 Formalization
In this section, we formalize our specification language, ESL. We begin with the definition of Deductive Normal Form (DeNF), which serves as a foundation for the remainder of this paper.
Definition 1.
Given a propositional formula $\psi_{1}\Rightarrow\psi_{2}$, if $\psi_{1}$ is a disjunctive norm form [48] and $\psi_{2}$ is a conjunctive norm form [48], then we call $\psi_{1}\Rightarrow\psi_{2}$ is a deductive normal form (DeNF).
Definition 2.
An ESL rule is a DeNF where each variable is substituted by a predicate.
An ESL specification ${\mathcal{E}}$ comprises three components: a variable set ${\mathcal{V}}$, a predicate set ${\mathcal{P}}$, and an ESL rule set ${\mathcal{R}}_{\mathcal{E}}$. For each predicate $f\in{\mathcal{P}}$, an associated natural language description is required. For instance, the following ESL specification encodes the regulation aforementioned in Section 3.1:
{minted}
[fontsize=,bgcolor=backgroundcolor,escapeinside=||]python "Variables": "x", "Predicates": "ChewGum(x) := a person x consumes or attempts to consume chewing gum or bublle gum.", "InRailway(x) := a person x is in some part of railway premises", "Rules": "InRailway(x) => not ChewGum(x)"
[…]
4 Methodologies of RvLLM
As illustrated in Figure 1, RvLLM conducts runtime verification through four stages: interpretation abstraction, rule normalization, forward chaining, and query generation. In this section, we detail the methodology of each stage and demonstrate its application using an example given in Figure 2.
4.1 Interpretation abstraction
Given a context and LLM outputs (the union denoted as $\sigma$) as the domain of discourse $\mathsf{D}$ and an ESL specification ${\mathcal{E}}=\langle\mathcal{V},\mathcal{P},\mathcal{R}_{\mathcal{E}}\rangle$, we first leverage a perception agent to obtain all possible objects from $\mathsf{D}$ and propositionalize as much as possible of the predicates defined in $\sigma$. We call this the perception process. Then, for each rule $r\in\mathcal{R}_{\mathcal{E}}$, we bind the variables in each predicate to all possible objects in a search-and-replace operation, to generate the propositional formulae set. Note that, given a variable binding for a rule $r\in\mathcal{R}_{\mathcal{E}}$, the interpretation has two different results: i) all predicates in the LHS of $r$ are propositionalized successfully with assigned truth value under the binding, which we call a complete binding in this work, or ii) only part of predicates are propositionalized due to the insufficient information from the domain of discourse, which we call a partial binding.
Given this distinction, we introduce two levels of interpretation in this work: Level-1 and Level-2. Specifically, Level-1 interpretation disregards all the partial variable binding and operates only on the complete variable bindings. Given a partial binding, Level-2 interpretation, in contrast, employs a perception agent to instantiate all the variables that are not assigned with an object in the binding, in a way such that the resultant proposition returns ${\sf True}$ given the domain of discourse.
Example 1.
Consider the example in Figure 2, which illustrates our approach using a simple numerical comparison question with a specification encoding a basic algebraic property of multiplication, the interpretation abstraction process first gets all possible objects and propositionalizes the predicate ${\sf IsGreater}$ as follows: $\{p_{0}={\sf IsGreater}(o_{1},o_{2})={\sf True},p_{1}={\sf IsGreater}(o_{2},o_{1})={\sf False}\}$ with $o_{1}=15.2$ and $o_{2}=15.12$. Then, it can obtain all possible variable binding $\{\flat_{1},\flat_{2}\}$ for the rule $\psi$ where $\flat_{1}=\{x\mapsto o_{1},y\mapsto o_{2}\}$ and $\flat_{2}\{x\mapsto o_{2},y\mapsto o_{1}\}$. $\flat_{1}$ results in a formula $p_{0}\wedge{\sf IsGreater}(z,0)\Rightarrow{\sf IsGreater}(15.2*z)$ for $\psi$, which is an partial binding. Similarly, $\flat_{2}$ is also partial. Consequently, for the Level-1 interpretation, no propositional formula is successfully generated, and the checking procedure exits with no inconsistency detected. While for the Level-2 interpretation, it successfully instantiates the variable $z$ for the partial biddings $\flat_{1}$ and $\flat_{2}$ with an object $o_{3}=10$, and obtain an additional proposition $p_{2}={\sf IsGreater}(o_{3},0)={\sf True}$. Then, the updated bindings are: $\flat_{1}^{\prime}=\{x\mapsto o_{1},y\mapsto o_{2},z\mapsto o_{3}\},\flat_{2}^{\prime}=\{x\mapsto o_{2},y\mapsto o_{1},z\mapsto o_{3}\}$. By applying $\flat_{1}^{\prime}$, we can obtain a new proposition $p_{3}={\sf IsGreater}(o_{1}*o_{3},o_{2}*o_{3})={\sf IsGreater}(15.2*10,15.12*10)$ but with unknown truth value. Similarly, we obtain $p_{4}={\sf IsGreater}(o_{2}*o_{3},o_{1}*o_{3})={\sf IsGreater}(15.12*10,15.2*10)={\sf Unknown}$ by $\flat_{2}^{\prime}$. Finally, we generate two propositional formulae: $\psi_{1}=p_{0}\wedge p_{2}\Rightarrow p_{3}$ via $\flat_{1}^{\prime}$ and $\psi_{2}=p_{1}\wedge p_{2}\Rightarrow p_{4}$ via $\flat_{2}^{\prime}$ for the following component.
Figure 2: Runtime verification of GPT 4.1 nano by RvLLM for a number comparison task.
4.2 Rule-like propositional formula transformation
Given a DeNF formula $\psi=D_{1}\vee D_{2}\vee\cdots\vee D_{m}\Rightarrow C_{1}\wedge C_{2}\wedge\cdots\wedge C_{n}$, where each $D_{i}$ is a conjunction of literals $l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}$ and each $C_{j}$ is a disjunction of literals $l_{j1}^{C}\vee l^{C}_{j2}\vee\cdots\vee l^{C}_{jn_{j}}$, the formula transformation module is designed to reformulate $\psi$ into a set of propositional formulae $\Gamma_{\psi}$ in a rule-like form, i.e., an implication, such that the LHS of each formula in $\Gamma_{\psi}$ is a literal conjunction and the RHS is a single literal.
To achieve this, we first transform $\psi$ by a formula set $\Gamma_{\psi}^{\prime}=\{D_{i}\Rightarrow C_{j}\mid i\in[m],j\in[n]\}$. Then, for each single formula $\psi_{i,j}=D_{i}\Rightarrow C_{j}=l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}\Rightarrow l_{j1}^{C}\vee l^{C}_{j2}\vee\cdots\vee l^{C}_{jn_{j}}$, we reformulate it into the following set of implicant formula:
$\Gamma_{\psi_{i,j}}^{\prime}=\{l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}\wedge(\bigwedge_{k\in[n_{j}]\wedge k\neq t}\neg l^{C}_{jk})\Rightarrow l^{C}_{jt}\mid t\in[n_{j}]\}$
Finally, given a DeNF $\psi$, we obtain the corresponding reformulated rule-like propositional formula set $\Gamma_{\psi}=\bigcup_{i\in[m],j\in[n]}\Gamma_{\psi_{i,j}}^{\prime}$, and given a DeNF set ${\mathcal{R}}$, we use $\Gamma_{\mathcal{R}}$ to denote the union of all rule-like form formulae transformed by each formula in ${\mathcal{R}}$, i.e., $\Gamma_{\mathcal{R}}=\bigcup_{\psi\in{\mathcal{R}}}\Gamma_{\psi}$. In the remainder of this paper, we use $\texttt{Lit}_{\mathcal{R}}$ (resp. $\texttt{Lit}_{\psi}$) to denote the set of all the literals that appeared in ${\mathcal{R}}$ (resp. $\psi$). The rules $\psi_{1}$ and $\psi_{2}$ obtained in Example 1 are already expressed in the rule-like form.
Example 2.
Consider a DeNF $\psi=(a_{1}\wedge b_{1})\vee a_{2}\Rightarrow(c_{1}\vee d_{1})\wedge c_{2}$. We first transform $\psi$ into an equivalent formula set $\Gamma_{\psi}^{\prime}=\{\psi_{1,1},\psi_{1,2},\psi_{2,1},\psi_{2,2}\}$, where $\psi_{1,1}=a_{1}\wedge b_{1}\Rightarrow c_{1}\vee d_{1}$, $\psi_{1,2}=a_{1}\wedge b_{1}\Rightarrow c_{2}$, $\psi_{2,1}=a_{2}\Rightarrow c_{1}\vee d_{1}$, and $\psi_{2,2}=a_{2}\Rightarrow c_{2}$. For each single formula in $\Gamma_{\psi}^{\prime}$, we then reformulate it as follows: $\Gamma_{\psi_{1,1}}^{\prime}=\{a_{1}\wedge b_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge b_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{1,2}}^{\prime}=\{a_{1}\wedge b_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,1}}^{\prime}=\{a_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,2}}^{\prime}=\{a_{2}\wedge c_{2}\}$, and finally obtain the set of rule-like propsitional formulas as $\Gamma_{\psi}=\Gamma_{\psi_{1,1}}^{\prime}\cup\Gamma_{\psi_{1,2}}^{\prime}\cup\Gamma_{\psi_{2,1}}^{\prime}\cup\Gamma_{\psi_{2,2}}^{\prime}$.
4.3 Forward chaining
[…]
To achieve it, given a rule set $\Gamma_{\mathcal{R}}$ and an initially defined literals Lit with truth values, we first construct an initial graph $G$ through the following steps:
  • Step 1 (Literal Nodes): Create a node for each literal that appears in the rule set $\Gamma_{\mathcal{R}}$.
  • Step 2 (LHS Node): Create an LHS node for each rule-like propositional formula in $\Gamma_{\mathcal{R}}$.
  • Step 3 (Edges): For each formula in $\Gamma_{\mathcal{R}}$, such as $l_{1}\wedge\cdots\wedge l_{n}\Rightarrow l_{n+1}\in\Gamma_{\mathcal{R}}$, add a directed edge from each literal node $l_{i}$ ($i\in[n]$) to the corresponding LHS node, and add a directed edge from this LHS node to the literal node $l_{n+1}$.
Next, according to the initially defined literals, we mark all the literal nodes valued ${\sf True}$ as $\texttt{Lit}^{\uparrow}$ and mark all the literal nodes valued ${\sf False}$ as $\texttt{Lit}^{\downarrow}$, based on the truth value of literals from Lit. Then, we perform forward chaining procedure on the graph $G$ and update $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\downarrow}$ as follows until no more update can be done or an inconsistency is detected by $\texttt{Lit}^{\uparrow}\cap\texttt{Lit}^{\downarrow}\neq\emptyset$:
  • Update of $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\downarrow}$: For each LHS node encoding the rule-like formula (or implication) $l_{1}\wedge\cdots\wedge l_{n}\Rightarrow l_{n+1}$, if $\{l_{1},\ldots,l_{n}\}\subseteq\texttt{Lit}^{\uparrow}$, then we have $\texttt{Lit}^{\uparrow}=\texttt{Lit}^{\uparrow}\cup l_{n+1}$ and $\texttt{Lit}^{\downarrow}=\texttt{Lit}^{\downarrow}\cup\neg l_{n+1}$.
Example 3.
We next illustrate the graph construction and the corresponding FC procedure for the rule set $\Gamma_{\psi}=\{\psi_{1},\psi_{2}\}$ under the Level-2 interpretation given in Example 1. We first obtain all the literals as $\texttt{Lit}_{\mathcal{R}}=\{p_{0},p_{1},p_{2},p_{3},p_{4}\}$ with $\texttt{Lit}^{\uparrow}=\{p_{0},p_{2}\}$ and $\texttt{Lit}^{\downarrow}=\{p_{1}\}$, construct the literal nodes correspondingly, and construct the LHS node set as $\{p_{0}\&p_{2},p_{1}\&p_{2}\}$. Then, directed edges are added as follows: $p_{0}\rightarrow p_{0}\&p_{2}$, $p_{2}\rightarrow p_{0}\&p_{2}$, $p_{0}\&p_{2}\rightarrow p_{3}$, $p_{1}\rightarrow p_{1}\&p_{2}$, $p_{2}\rightarrow p_{1}\&p_{2}$, $p_{1}\&p_{2}\rightarrow p_{4}$, resulting an intial graph. Next, we check the logic consistency. For the LHS node $p_{0}\&p_{2}$, since $\{p_{0},p_{2}\}\subseteq\texttt{Lit}^{\uparrow}$, we can update $\texttt{Lit}^{\uparrow}$ as $\texttt{Lit}^{\uparrow}=\texttt{Lit}^{\uparrow}\cup p_{3}=\{p_{0},p_{2},p_{3}\}$ and $\texttt{Lit}^{\downarrow}=\texttt{Lit}^{\downarrow}\cup\neg p_{3}=\{p_{1},\neg p_{3}\}$, where “$p_{3}={\sf True}$” is the new knowledge derived. For the LHS node $p_{1}\&p_{2}$, since $p_{1}\notin\texttt{Lit}^{\uparrow}$, there is no update happens to $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\uparrow}$; consequently, no new knowledge is inferred.
4.4 Query generation
Once RvLLM obtains the newly inferred knowledge, the query generation module generates a concrete query to the target LLM related to the knowledge, and requires the LLM to analyze its truth value. For the inferred knowledge $p_{3}={\sf IsGreater}(151.2,152)={\sf True}$ in Example 3, we generate the corresponding query as “Is 151.2 greater than 152?”, as shown in Figure 2. Since the target LLM returns ${\sf False}$, an inconsistency is detected, indicating a diagnostic error in the target LLM.
5 Experiments
To evaluate the effectiveness of RvLLM, we apply it to the runtime verification of LLMs across three representative tasks that require domain-specific knowledge: violation detection against Singapore Rapid Transit Systems Act [1], numerical comparison, and inequality solving. For a comprehensive evaluation, we include both SOTA and non-SOTA LLMs as our benchmark, including Qwen 2.5 (max, plus, turbo, 7B, 14B, 32B, 72B), GPT (4.1, 4.1 mini, 4.1 nano), Gemini 2.0 (Flash, Flash Lite), and DeepSeek-V3. All experiments are conducted on a machine with an Intel (R) Xeon(R) w7-2475X processor. A dedicated guideline for designing interpretation abstraction prompts for the perception agent is provided in the appendices, along with additional experimental details and descriptions of the datasets and specifications used throughout the experiments.
5.1 Case study 1: violation detection against Singapore Rapid Transit Systems Act
For this case study, we apply RvLLM under the Level-1 interpretation to evaluate the LLM’s responses in detecting violations within contextual scenarios against the Singapore Rapid Transit Systems Act. In this capacity, RvLLM can be regarded as a complementary mechanism to enhance the LLM’s ability in detecting the law violations in this study. We conduct experiments using contextual scenarios derived from [53], consisting of 304 cases in total–280 labeled as involving violations (unsafe) and 24 as not (safe). To incorporate relevant domain knowledge, we carefully encode 31 of the 53 regulations from Singapore Rapid Transit System Act into ESL specifications. It took a research scientist three days to interpret the law and develop corresponding specifications–a one-time effort applicable for verifying any LLM application against these regulations. In this study, the same model serves both as the target LLM and as the perception agent.
Results. Table 1 reports the violation detection results by various LLMs, with and without our framework as a complementary analysis method, where TPR and TNR denote the true positive rate and true negative rate for the violation detection, i.e., a violation is a positive example. The last column gives the average execution time for each scenario analysis. In the standalone LLM setting, a true positive is defined as a case where the model successfully identifies a violation in an unsafe scenario, and a true negative corresponds to a case where the model confirms the absence of violations in a safe scenario. In the combined LLM+RvLLM setting, a true positive is recorded if either the LLM or RvLLM detects a violation in an unsafe scenario, whereas a true negative is defined as a case that both the LLM and RvLLM determine the absence of any violation in a safe scenario. The results show that integrating RvLLM significantly enhances the TPR, improving the target LLM’s capability to identify violations. However, we also observe that the integration of RvLLM can reduce the TNR in some models. This decline is primarily due to inaccuracies or incompleteness in the interpretation procedure by the perception agent. We also find that, among all evaluated models, DeepSeek-V3 achieves the highest TPR in both standalone and integrated settings, demonstrating superior detection capabilities in this domain-specific task and good performance on language understanding.
Table 1: Violation detection results of LLMs with and without RvLLM against the Singapore Rapid Transit Systems Act [1].
LLM LLM + RvLLM Performance of RvLLM
Target LLMs TPR TNR TPR TNR TPR TNR Time (s)
Qwen max [56] 56.2% 91.3% 86.1% 91.3% 29.9% $\uparrow$ 0 3.00
Qwen plus [56] 58.4% 1 81.9% 1 23.5% $\uparrow$ 0 5.32
Qwen turbo [56] 39.3% 1 75% 1 35.7% $\uparrow$ 0 1.64
Qwen 2.5 (7B) [56] 16.0% 1 61.2% 87.0% 45.2% $\uparrow$ 13% $\downarrow$ 2.28
Qwen 2.5 (14B) [56] 38.8% 95.7% 80.4% 87.0% 41.6% $\uparrow$ 8.7% $\downarrow$ 3.11
Qwen 2.5 (32B) [56] 56.6% 95.7% 87.5% 95.7% 30.9% $\uparrow$ 0 2.55
Qwen 2.5 (72B) [56] 57.3% 1 80.4% 95.7% 23.1% $\uparrow$ 4.3% $\downarrow$ 2.57
GPT 4.1 [3] 57.7% 95.7% 81.1% 91.3% 23.4% $\uparrow$ 4.3% $\downarrow$ 3.11
GPT 4.1 mini [3] 39.1% 95.7% 65.1% 95.7% 26% $\uparrow$ 0 3.94
GPT 4.1 nano [3] 11.8% 1 30.0% 1 18.2% $\uparrow$ 0 1.82
Gemini 2.0 Flash [21] 37.0% 1 87.2% 82.6% 50.2% $\uparrow$ 17.4% $\downarrow$ 2.69
Gemini 2.0 Flash Lite [21] 55.1% 95.7% 79.7% 95.7% 24.6% $\uparrow$ 0 1.72
DeepSeek-V3 [39] 78.3% 87.0% 94.0% 82.6% 15.7% $\uparrow$ 4.4% $\downarrow$ 15.84
5.2 Case study 2: numerical comparison problems
For this case study, we apply RvLLM under Level-2 interpretation to verify the numerical comparison results produced by various LLMs. It is important to note that RvLLM is inherently designed without mathematical solving capabilities and relies solely on logical inference. Without loss of generality, we randomly generate 100 questions following the guideline for increasing the likelihood of incorrect comparison results by LLMs. The specification file used here is the one shown in Figure 2. Again, we use the same model to serve both the target LLM and the perception agent.
Table 2: Runtime verification results of various LLMs on numerical comparison tasks.
Target LLMs Con. Incon. Fail Time (s)
Qwen max [56] 93 7 0 5.03
Qwen plus [56] 99 1 0 7.01
Qwen turbo [56] 94 6(2) 0 2.93
Qwen 2.5 (7B) [56] 0 0 100 N/A
Qwen 2.5 (14B) [56] 90 10(7) 0 5.09
Qwen 2.5 (32B) [56] 98 2 0 4.98
Qwen 2.5 (72B) [56] 96 4 0 6.31
GPT 4.1 [3] 98 2 0 4.84
GPT 4.1 mini [3] 92(1) 8 0 5.37
GPT 4.1 nano [3] 90(3) 10 0 3.73
Gemini 2.0 Flash [21] 93 7 0 4.94
Gemini 2.0 Flash Lite [21] 52 44 4 4.35
DeepSeek-V3 [39] 69 31(5) 0 30.87
Results. Table 2 gives the verification results by RvLLM on numerical comparison tasks across various LLMs. Columns Con. and Incon. report the number of cases where RvLLM detects no inconsistencies and where at least one inconsistency is detected, following the process in Figure 2. The values in parentheses in Column Con. represent cases where the target LLM produces an erroneous comparison result on the original comparison question, yet RvLLM fails to identify any inconsistencies–typically due to incomplete interpretation by the perception agents. In contrast, the values in parentheses in Column Incon. corresponds to cases where the target LLM returns a correct comparison result on the original question, but RvLLM indeed detects some inconsistencies. This usually arises from two main reasons: i) incomplete interpretation, and ii) the target LLM producing an incorrect diagnosis on newly-inferred knowledge. Overall, RvLLM effectively identifies erroneous diagnoses by the target LLM in most cases, with exceptions observed in only 3 cases for GPT 4.1 nano and mini. Among all tested models, Gemini 2.0 Flash Lite and DeepSeek-V3 show the weakest diagnostic performance, and Qwen 2.5 (7B) has the worst language understanding capabilities, failing all interpretation abstraction. Given its strong performance in the previous violation detection task, we conjecture that DeepSeek-V3 is better suited for language understanding tasks rather than those requiring rigorous logical reasoning.
5.3 Case study 3: inequality solving problems
For this case study, we employ RvLLM under Level-1 interpretation to verify the step-by-step reasoning process of LLMs in inequality solving tasks. We compile a dataset of 40 inequality questions sourced from A-Level H2 Mathematics examination papers [2] and carefully design three specifications to serve as domain knowledge that RvLLM uses for verification. These specifications consist of basic algebraic inequality properties taught at the junior college level, targeting three common LLM reasoning errors in inequality solving: incorrect factorization, flawed interval analysis, and omission of endpoint or critical point checks. Given the higher interpretation complexity in this case study and the strong language processing capabilities of DeepSeek-V3, we employ it as the perception agent for all LLMs evaluated and Qwen 2.5 (32B) as an alternative for comparison.
Results. Given the strong performance of RvLLM in detecting true positives, we focus the evaluation on the questions that the target LLMs answered incorrectly. The results are summarized in Table 3. Column 2 lists the total number of questions for which the target LLM produces incorrect solutions. Columns labeled DS-V3 and QW-32B (spanning Columns 3 to 11) show the number of cases where RvLLM successfully identifies inconsistencies based on the corresponding specifications using DeepSeek-V3 and Qwen 2.5 (32B) as the perception agent, respectively. Values in parentheses indicate false positives–cases where the detected inconsistency does not stem from an actual reasoning error but from inaccuracies of the perception agent during the interpretation abstraction process. Column GT gives the ground truth number of incorrect responses attributable to the rule specifications. It is important to note that we define only three specifications as domain knowledge for inequality-solving tasks, which is insufficient to comprehensively cover all reasoning patterns required for such problems. As a result, it is expected that RvLLM fails to detect a portion of the errors due to the incompleteness of the specifications. The final two columns give the true positive rate of RvLLM with three specifications. We find that RvLLM achieves a TPR of up to 52.2% when using DeepSeek-V3 as the perception agent. This result highlights not only the potential of RvLLM in detecting reasoning errors (even with incomplete domain knowledge), but also underscores the importance of employing a powerful perception agent to enhance the overall performance. Note that, although the TPR remains below 50% for most of the models, it can reach significantly higher values when considering only the ground truth cases listed in Column GT, indicating the potential for performance improvement if more domain knowledge with respect to the inequality solving is incorporated.
Table 3: Verification results by RvLLM in inequality solving tasks with three specifications, defining rules related to factorization, interval analysis, and endpoints/critical points checking.
Factorization Error Interval Error Endpoints Error TPR
Target LLMs Incorrect DS-V3 QW-32B GT DS-V3 QW-32B GT DS-V3 QW-32B GT DS-V3 QW-32B
Qwen max [56] 12 1 1 1 1 0 1 2 2 3 33.3% 25%
Qwen plus [56] 10 1 1 2 0 0 0 0 1 2 10% 20%
Qwen turbo [56] 22 0 0 1 1 0 1 3 2 4 18.2% 9.1%
Qwen 2.5 (7B) [56] 23 2 1 3 7 2 7 3 2 5 52.2% 21.7%
Qwen 2.5 (14B) [56] 18 0 0 1 3 2(1) 3 3(1) 2 3 27.8% 16.7%
Qwen 2.5 (32B) [56] 16 1 0 1 3 2 4 4 2 5 50% 25%
Qwen 2.5 (72B) [56] 15 1 0 1 4 3 4 0 0 1 33.3% 20
GPT 4.1 [3] 5 1 1(1) 1 0 0 0 0 0 0 20% 0
GPT 4.1 mini [3] 2 0 0 0 0 0 0 1 0 1 50% 0
GPT 4.1 nano [3] 10 0 0 0 1 0 1 1 0 1 20% 0
Gemini 2.0 Flash [21] 5 0 0 0 1 0 1 0 0 0 20% 0
Gemini 2.0 Flash Lite [21] 6 1 1 2 0 0 0 0 0 1 16.7% 16.7%
6 Conclusion
In this work, we present the first runtime verification framework RvLLM for LLMs, which allows the incorporation of domain knowledge. To support this, we design a general specification language, ESL, which enables domain experts to formally encode constraints in a lightweight yet expressive manner, supporting the rigorous verification of LLM behavior. We implement the proposed framework as a tool and evaluate its effectiveness on three representative tasks where domain knowledge plays a critical role. For each task, we develop tailored specifications grounded in relevant domain expertise. Experimental results indicate that RvLLM effectively leverages expert-defined domain knowledge to enable precise runtime verification of LLMs. However, despite its good performance, the current framework suffers from limited expressiveness, restricting its ability to encode more complex domain knowledge and behavioral constraints for LLMs. Furthermore, the reasoning mechanism remains limited to a single strategy, as embodied in the forward chaining module. Future work will focus on enriching the specification language with additional operators to enhance its expressiveness and investigating the integration of more reasoning strategies. This is intended to enable a more nuanced and comprehensive inference process, thereby enhancing overall verification capabilities.

### v3

@@
-leftmargin=1em, rightmargin=1em, indentfirst=false
+RvLLM: LLM Runtime Verification with Domain Knowledge
 Abstract
@@
 In this work, we advance misbehavior detection by incorporating domain knowledge.
-The core idea is to design a general specification language that enables domain experts to customize domain-specific predicates in a lightweight and intuitive manner, supporting later runtime verification of LLM outputs.
+The core idea is to design a general specification language that enables domain experts to customize domain-specific constraints in a lightweight and intuitive manner, supporting later runtime monitoring of LLM outputs.
 To achieve this, we design a novel specification language ESL and introduce a runtime verification framework RvLLM to validate LLM output against domain-specific constraints defined in ESL.
 RvLLM operates in two main stages: interpretation and reasoning.
-During interpretation, it derives interpretations of the specification based on context, which then guide the reasoning process to identify inconsistencies.
+During interpretation, it derives interpretations of the specification based on the context, which then guide the reasoning process to identify inconsistencies.
 When new knowledge is derived, RvLLM issues a follow-up query to the LLM to further verify the consistency.
 We evaluate RvLLM on three representative tasks: violation detection against Singapore Rapid Transit Systems Act, numerical comparison, and inequality solving.
-Experimental results demonstrate that RvLLM effectively detects erroneous outputs across various LLMs in a lightweight and flexible manner.
-The results reveal that despite their impressive capabilities, LLMs remain prone to low-level errors due to limited interpretability and a lack of formal guarantees during inference, and our framework offers a potential long-term solution by leveraging expert domain knowledge to rigorously and efficiently verify LLM outputs.
+Experimental results show that RvLLM effectively detects erroneous outputs across various LLMs in a lightweight and flexible manner.
+The results reveal that despite their impressive capabilities, LLMs remain prone to low-level errors due to a lack of formal guarantees during inference, and our framework offers a potential long-term solution by leveraging expert domain knowledge to rigorously and efficiently verify LLM outputs.
 1 Introduction
-Unlike rule-based systems [47] operating on predefined and deterministic rules, large language models (LLMs) [3, 21, 56, 39] learn data representation and processing automatically from the training datasets, achieving human-like or even superhuman performance, and have driven significant advancements in various practical applications [11, 30, 50, 40, 22].
-However, their non-deterministic and unpredictable nature sometimes leads to inconsistent or erroneous outputs [5, 45], posing significant risks in safety-critical or knowledge-intensive domains.
-Recent studies [55, 6] have shown that such misbehavior in LLMs cannot be fully eliminated, underscoring the urgent need for dedicated verification and validation methodologies to enhance the reliability of LLM-generated outputs.
-LLM testing [59, 25, 28, 60] primarily aims to establish comprehensive benchmarks to evaluate the overall model performance against domain-agnostic criteria–such as accuracy, coherence, and fairness–in alignment with the intended application.
+Unlike rule-based systems [russell2016artificial] operating on predefined and deterministic rules, large language models (LLMs) [achiam2023gpt, geminiteam2024gemini, yang2024qwen2, liu2024deepseek] learn data representation and processing automatically from the training datasets, achieving human-like or even superhuman performance, and have driven significant advancements in various practical applications [cascella2023evaluating, kasneci2023chatgpt, thirunavukarasu2023large, liu2023your, guha2023legalbench].
+However, their non-deterministic and unpredictable nature sometimes leads to inconsistent or erroneous outputs [lawBBC, suicideBBC], posing significant risks in safety-critical or knowledge-intensive domains.
+Recent studies [xu2024hallucination, banerjee2024llms] have shown that such misbehavior in LLMs cannot be fully eliminated, underscoring the urgent need for dedicated verification and validation methodologies to enhance the reliability of LLM-generated outputs.
+LLM testing [zhong2024agieval, HendrycksBBZMSS21, huang2024c, zhou2023don] primarily aims to establish comprehensive benchmarks to evaluate the overall model performance against domain-agnostic criteria–such as accuracy, coherence, and fairness–in alignment with the intended application.
 While these approaches effectively assess general behavior and reveal edge cases that may provoke unexpected responses, they are limited to predefined benchmarks and lack the specificity needed to address domain-specific assessment needs.
 LLM verification, instead, may serve as a complementary mechanism to LLM testing by providing formal guarantees on model behavior.
-Despite substantial progress over the past decade in neural network verification [9, 17, 27, 31, 54, 44, 57], existing methods exhibit notable limitations: they are primarily tailored to simpler model architectures–such as deep or convolutional networks–and classification tasks, struggling to scale to the complexity of modern AI models and their diverse functionalities.
-Furthermore, these approaches target at general properties such as robustness and fairness, making them ill-suited for domain-specific properties.
+Despite substantial progress over the past decade in neural network verification [BunelLTTKK20, ElboherGK20, HKWW17, KBDJK17, betaCrown, narodytska2018verifying, BDD4BNN], existing methods exhibit notable limitations: they are primarily tailored to simpler model architectures–such as deep or convolutional networks–and classification tasks, struggling to scale to the complexity of modern AI models and their diverse functionalities.
+Furthermore, these approaches target general properties such as robustness and fairness, making them ill-suited for domain-specific properties.
 Therefore, developing a dedicated verification paradigm tailored to LLMs is crucial for ensuring certified and reliable outputs.
-On the other hand, the decline of rule-based expert systems [23, 29, 20] in the late 20th century can be attributed to their inability to handle incomplete information and poor scalability in real-world complexity.
-Defining comprehensive rule sets for open domains has been proven infeasible [47, 26].
+On the other hand, the decline of rule-based expert systems [hayes1983building, jackson1986introduction, forsyth1984expert] in the late 20th century can be attributed to their inability to handle incomplete information and poor scalability in real-world complexity.
+Defining comprehensive rule sets for open domains has been proven infeasible [russell2016artificial, hernandez2017measure].
 We argue that such rules can only be statistically approximated–a capability exemplified by LLMs.
@@
 This highlights the need for verifying explicitly defined rules in a lightweight and incremental manner for practical deployment and maintainability.
-Figure 1: An overview of RvLLM.
-Given the LLM’s context and outputs, RvLLM first extracts relevant propositions and translates the rules from the specification file into a normalized set of propositional logic formulae.
-These formulae are then processed via forward chaining to detect inconsistencies or infer new knowledge, which is then used to generate follow-up queries and assess consistency across successive LLM responses.
 This paper.
-Building on these foundations, we propose runtime verification with domain knowledge as a sustainable solution to ensuring reliable LLM behavior.
-Our work is motivated by two key insights: i) Runtime verification offers a lightweight yet rigorous means to bridge testing and formal verification by assessing system behavior against formal properties during execution; ii) Existing work primarily focuses on general misbehavior detection, and often overlooks the critical role of domain-specific expertise.
-Integrating such knowledge is essential for handling edge cases and improving LLM reliability in specialized tasks, as domain knowledge can significantly enhance model performance.
+Building on these foundations, we propose runtime verification with domain knowledge as a sustainable solution to ensuring reliable LLM behavior, motivated by two key insights: i) Runtime verification offers a lightweight yet rigorous means to bridge testing and formal verification by assessing system behavior against formal properties during execution; ii) Existing work primarily focuses on generic misbehavior detection while often overlooking the critical role of domain-specific expertise.
+Integrating such knowledge is crucial for handling edge cases and enhancing LLM reliability in specialized tasks, where domain knowledge can significantly improve performance.
 To achieve this, we introduce RvLLM, an innovative runtime verification framework tailored to LLMs to ensure reliable outputs, particularly against axiomatic domain specifications.
-The idea is to automatically check whether the LLM-generated responses adhere to domain-specific properties expressed in a simple, adaptable specification language.
+The idea is to automatically check whether the LLM-generated responses adhere to domain-specific constraints expressed in a simple, adaptable specification language.
 This simplicity and flexibility empower domain experts to encode their knowledge and formally define constraints that capture the expected behavior of LLMs in specialized applications.
@@
 Once new knowledge is inferred, the query generation module issues targeted queries, evaluates the LLM responses, and detects inconsistencies when outputs contradict the previously inferred knowledge.
+Figure 1: An overview of RvLLM.
+Given an LLM’s context and outputs, RvLLM first extracts relevant propositions and translates the ESL specification into normalized propositional logic formulae.
+These are then processed through forward chaining to detect inconsistencies and infer new knowledge, which subsequently guides follow-up queries and consistency checks across successive LLM responses.
 Experimental results.
-We evaluate the effectiveness of RvLLM through experiments on three representative tasks: violation detection against Singapore Rapid Transit Systems Act [1], numerical comparison, and inequality solving, using a diverse set of LLMs.
+We evaluate the effectiveness of RvLLM through experiments on three representative tasks: violation detection against Singapore Rapid Transit Systems Act [SGMRT], numerical comparison, and inequality solving, using a diverse set of LLMs.
 The experimental results show that RvLLM significantly enhances the reliability of LLM output in these domain-specific tasks.
-In the violation detection task, employing RvLLM as a complementary mechanism increases the true positive rate, or TPR, by 15.7% to 50.2% across various models.
+In the violation detection task, employing RvLLM as a complementary mechanism increases the true positive rate (TPR) by 15.7% to 50.2% across various models.
 In the numerical comparison task, RvLLM detects nearly all errors in the LLM responses across 100 randomly generated questions.
@@
 These specifications are primarily grounded in fundamental algebraic inequality properties typically covered at the junior college level.
-Notably, despite the incomplete domain knowledge in the inequality solving task, RvLLM still achieves a TPR of 52.2% in detecting erroneous solutions.
+Notably, despite the incomplete domain knowledge in the inequality solving task, RvLLM still achieves a TPR of 50% in detecting erroneous solutions.
 We argue that such a runtime verification approach can function as a critical safeguard for LLMs, particularly in rule-based or reasoning-intensive domains where adherence to specific constraints is essential.
-2 Related work
-LLM testing.
-Research in this area aims to develop comprehensive benchmarks and systematic evaluation frameworks to assess LLM across general attributes such as robustness, fairness, and factual consistency [41, 37, 52].
-Robustness studies examine model resilience to adversarial input, paraphrasing, or noise [10, 51, 58], while fairness evaluations assess bias with respect to sensitive attributes like gender or race [35, 58].
-Factual consistency detection, or hallucination detection, focuses on identifying instances in which LLMs produce plausible yet factually incorrect or unsupported claims [38, 34, 36].
-Although large-scale benchmarks like HELM [37] and BIG-bench [49] offer task suites and metrics for rigorous performance evaluation, spanning accuracy, calibration, and ethical considerations, a critical challenge persists, i.e., the dynamic and open-ended nature of real-world deployments complicates the creation of exhaustive and representative test suites.
-Runtime verification of LLMs.
-Runtime verification [7, 19] is a dynamic analysis technique that ensures certified system behavior during execution by monitoring traces against formal specifications.
-Owing to its lightweight nature, runtime verification has inspired multiple efforts to verify LLM-generated outputs at runtime [14, 8, 43, 12, 33, 42, 32].
-However, existing approaches center on open-domain settings, where specifications are often ambiguous and informal.
-For instance, SelfCheckGPT [43] and CHECKEMBED [8] operationalize specifications in terms of output stability.
-More recently, [13] proposes a fairness monitoring approach leveraging precisely defined properties in formal specifications.
-These specifications, informed by linear temporal logic [46] and its bounded metric variant [4], signal a shift towards formal methods for more accurate and dependable runtime verification of LLM behaviors.
-However, these works primarily focus on general properties and are not well suited to domain-specific constraints for specialized tasks.
+We further compare RvLLM with existing runtime hallucination-detection methods.
+Results show that RvLLM demonstrates superior performance by maintaining both high true positive and negative rates, whereas existing methods often exhibit a non-negligible trade-off between these metrics.
 2 ESL: a simple way of specifying domain-specific properties
@@
 The core objective is to strike a balance between user-friendliness and expressiveness.
-We illustrate this with the following regulation from Singapore Rapid Transit System Act [1]: {quoting} “No person shall consume or attempt to consume any chewing gum or bubble gum while in or upon any part of the railway premises.”
+We illustrate this with the following regulation from Singapore Rapid Transit System Act [SGMRT]:
+  • “No person shall consume or attempt to consume any chewing gum or bubble gum while in or upon any part of the railway premises.”
 Accessibility for domain experts.
 The primary goal of this specification language is to empower domain experts, even those with limited programming or formal methods experience, to effectively define and enforce constraints on the LLM’s behavior.
-To balance expressiveness and user-friendliness, we base the specification rule on predicate logic [48, 18], restricted to prenex normal forms [48, 18] and consisting solely of universal quantifiers.
+To balance expressiveness and user-friendliness, we base the specification rule on predicate logic [shoenfield2018mathematical, enderton2001mathematical], restricted to prenex normal forms [shoenfield2018mathematical, enderton2001mathematical] and consisting solely of universal quantifiers.
 We further constrain its quantifier-free part to a rule-like form (or implication), yielding a simplified variant of predicate logic expressible as $\forall\vec{x}.(f(\vec{x})\Rightarrow g(\vec{x}))$ [^1] We use $\vec{x}$ instead of $x$ as there could be a sequence of variables..
-Specifically, both the left-hand side (LHS) and right-hand side (RHS) of the rule-like part are quantifier-free predicate formulas composed of predicates applied to terms that may include variables (e.g., $x$), constants (e.g., ‘bubble gum’), and functions of terms (e.g., $\text{sin}(x)$).
+Specifically, both the left-hand side (LHS) and right-hand side (RHS) of the rule-like part are quantifier-free predicate formulae composed of predicates applied to terms that may include variables (e.g., $x$), constants (e.g., “bubble gum”), and functions of terms (e.g., $\text{sin}(x)$).
 For example, ${\sf ChewGum}(x)$ is an atomic predicate with a single variable argument $x$, and ${\sf IsGreater}({\sf Square}(x),0)$ is also an atomic predicate with two arguments: the function term ${\sf Square}(x)$ and the constant term 0.
@@
 For example, a domain expert can define their specialized atomic predicates ${\sf ChewGum}(x)$ and ${\sf InRailway}(x)$ associated with a natural language description as follows:
-{quoting}
-[…]
 […]
@@
 Definition 1.
-Given a propositional formula $\psi_{1}\Rightarrow\psi_{2}$, if $\psi_{1}$ is a disjunctive norm form [48] and $\psi_{2}$ is a conjunctive norm form [48], then we call $\psi_{1}\Rightarrow\psi_{2}$ is a deductive normal form (DeNF).
+Given a propositional formula $\psi_{1}\Rightarrow\psi_{2}$, if $\psi_{1}$ is a disjunctive norm form [shoenfield2018mathematical] and $\psi_{2}$ is a conjunctive norm form [shoenfield2018mathematical], then we call $\psi_{1}\Rightarrow\psi_{2}$ is a deductive normal form (DeNF).
 Definition 2.
-An ESL rule is a DeNF where each variable is substituted by a predicate.
+An ESL rule is a DeNF where each proposition is substituted by a predicate.
 An ESL specification $\mathcal{E}$} comprises three components: a variable set $\mathcal{V}$}, a predicate set $\mathcal{P}$}, and an ESL rule set $\mathcal{R}$$_{\mathcal{E}}$.
@@
 For instance, the following ESL specification encodes the regulation aforementioned in Section 2.1:
-{minted}
-[fontsize=,bgcolor=backgroundcolor,escapeinside=||]python "Variables": "x", "Predicates": "ChewGum(x) := a person x consumes or attempts to consume chewing gum or bublle gum.", "InRailway(x) := a person x is in some part of railway premises", "Rules": "InRailway(x) => not ChewGum(x)"
-[…]
+"Variables": {"x"},
+"Predicates": {"ChewGum(x) := a person x consumes or attempts to consume chewing gum or bubble gum.",
+"InRailway(x) := a person x is in some part of railway premises"},
+"Rules": {"InRailway(x) => not ChewGum(x)"}
+Interpretation of ESL rules.
+An interpretation of an ESL rule $r$ assigns objects to variables (i.e., variable binding) in the predicate of $r$ and evaluates the truth value ($\sf True$}, $\sf False$}, or $\sf Unknown$}) of the resulting proposition after substitution.
+Note that we adopt the open-world assumption [drummond2006open] to better accommodate the open-ended nature of LLMs.
+In this work, interpretations are derived from the context and LLM outputs, serving as the domain of discourse [hein2003discrete].
+Now, consider the following contextual scenario:
+  • “In a crowded MRT train, Alex nervously chews gum to ease stress before an interview.
+Suddenly, the train jolts, and the gum flies out, landing on a stranger’s shirt.
+Awkward glances turn into laughter as apologies spill out, diffusing tension in the confined space.”
+The interpretation of the rule ${\sf InRailway}(x)\Rightarrow\neg{\sf ChewGum}(x)$ is $\{{\sf InRailway}(o_{1})={\sf True}$, ${\sf InRailway}(o_{2})={\sf True}$, ${\sf ChewGum}(o_{1})={\sf True}$, ${\sf ChewGum}(o_{2})={\sf Unknown}\}$, where $o_{1}=\text{`Alex'}$ and $o_{2}=\text{`stranger'}$.
 3 Methodologies of RvLLM
@@
 Consequently, for the Level-1 interpretation, no propositional formula is successfully generated, and the checking procedure exits with no inconsistency detected.
-While for the Level-2 interpretation, it successfully instantiates the variable $z$ for the partial biddings $\flat_{1}$ and $\flat_{2}$ with an object $o_{3}=10$, and obtain an additional proposition $p_{2}={\sf IsGreater}(o_{3},0)={\sf True}$.
+While for the Level-2 interpretation, it successfully instantiates the variable $z$ for the partial bindings $\flat_{1}$ and $\flat_{2}$ with an object $o_{3}=10$, and obtain an additional proposition $p_{2}={\sf IsGreater}(o_{3},0)={\sf True}$.
 Then, the updated bindings are: $\flat_{1}^{\prime}=\{x\mapsto o_{1},y\mapsto o_{2},z\mapsto o_{3}\},\flat_{2}^{\prime}=\{x\mapsto o_{2},y\mapsto o_{1},z\mapsto o_{3}\}$.
@@
 Given a DeNF formula $\psi=D_{1}\vee D_{2}\vee\cdots\vee D_{m}\Rightarrow C_{1}\wedge C_{2}\wedge\cdots\wedge C_{n}$, where each $D_{i}$ is a conjunction of literals $l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}$ and each $C_{j}$ is a disjunction of literals $l_{j1}^{C}\vee l^{C}_{j2}\vee\cdots\vee l^{C}_{jn_{j}}$, the formula transformation module is designed to reformulate $\psi$ into a set of propositional formulae $\Gamma_{\psi}$ in a rule-like form, i.e., an implication, such that the LHS of each formula in $\Gamma_{\psi}$ is a literal conjunction and the RHS is a single literal.
-To achieve this, we first transform $\psi$ by a formula set $\Gamma_{\psi}^{\prime}=\{D_{i}\Rightarrow C_{j}\mid i\in[m],j\in[n]\}$.
-Then, for each single formula $\psi_{i,j}=D_{i}\Rightarrow C_{j}=l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}\Rightarrow l_{j1}^{C}\vee l^{C}_{j2}\vee\cdots\vee l^{C}_{jn_{j}}$, we reformulate it into the following set of implicant formula:
+To achieve this, we first transform $\psi$ to a formula set $\Gamma_{\psi}^{\prime}=\{D_{i}\Rightarrow C_{j}\mid i\in[m],j\in[n]\}$.
+Then, for each single formula $\psi_{i,j}=D_{i}\Rightarrow C_{j}=l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}\Rightarrow l_{j1}^{C}\vee l^{C}_{j2}\vee\cdots\vee l^{C}_{jn_{j}}$, we reformulate it as the implicant formula set $\Gamma_{\psi_{i,j}}^{\prime}$, following the idea of unit propagation [zhang1996cient]:
 $\Gamma_{\psi_{i,j}}^{\prime}=\{l_{i1}^{D}\wedge l^{D}_{i2}\wedge\cdots\wedge l^{D}_{im_{i}}\wedge(\bigwedge_{k\in[n_{j}]\wedge k\neq t}\neg l^{C}_{jk})\Rightarrow l^{C}_{jt}\mid t\in[n_{j}]\}$
@@
 We first transform $\psi$ into an equivalent formula set $\Gamma_{\psi}^{\prime}=\{\psi_{1,1},\psi_{1,2},\psi_{2,1},\psi_{2,2}\}$, where $\psi_{1,1}=a_{1}\wedge b_{1}\Rightarrow c_{1}\vee d_{1}$, $\psi_{1,2}=a_{1}\wedge b_{1}\Rightarrow c_{2}$, $\psi_{2,1}=a_{2}\Rightarrow c_{1}\vee d_{1}$, and $\psi_{2,2}=a_{2}\Rightarrow c_{2}$.
-For each single formula in $\Gamma_{\psi}^{\prime}$, we then reformulate it as follows: $\Gamma_{\psi_{1,1}}^{\prime}=\{a_{1}\wedge b_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge b_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{1,2}}^{\prime}=\{a_{1}\wedge b_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,1}}^{\prime}=\{a_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,2}}^{\prime}=\{a_{2}\wedge c_{2}\}$, and finally obtain the set of rule-like propsitional formulas as $\Gamma_{\psi}=\Gamma_{\psi_{1,1}}^{\prime}\cup\Gamma_{\psi_{1,2}}^{\prime}\cup\Gamma_{\psi_{2,1}}^{\prime}\cup\Gamma_{\psi_{2,2}}^{\prime}$.
+For each single formula in $\Gamma_{\psi}^{\prime}$, we then reformulate it as follows: $\Gamma_{\psi_{1,1}}^{\prime}=\{a_{1}\wedge b_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge b_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{1,2}}^{\prime}=\{a_{1}\wedge b_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,1}}^{\prime}=\{a_{1}\wedge\neg c_{1}\Rightarrow d_{1},a_{1}\wedge\neg d_{1}\Rightarrow c_{1}\}$, $\Gamma_{\psi_{2,2}}^{\prime}=\{a_{2}\wedge c_{2}\}$, and finally obtain the set of rule-like propositional formulae as $\Gamma_{\psi}=\Gamma_{\psi_{1,1}}^{\prime}\cup\Gamma_{\psi_{1,2}}^{\prime}\cup\Gamma_{\psi_{2,1}}^{\prime}\cup\Gamma_{\psi_{2,2}}^{\prime}$.
 3.3 Forward chaining
@@
 To achieve it, given a rule set $\Gamma_{\mathcal{R}}$ and an initially defined literals Lit with truth values, we first construct an initial graph $G$ through the following steps:
+-1mm
   • Step 1 (Literal Nodes): Create a node for each literal that appears in the rule set $\Gamma_{\mathcal{R}}$.
@@
 Then, we perform forward chaining procedure on the graph $G$ and update $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\downarrow}$ as follows until no more update can be done or an inconsistency is detected by $\texttt{Lit}^{\uparrow}\cap\texttt{Lit}^{\downarrow}\neq\emptyset$:
+-1mm
   • Update of $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\downarrow}$: For each LHS node encoding the rule-like formula (or implication) $l_{1}\wedge\cdots\wedge l_{n}\Rightarrow l_{n+1}$, if $\{l_{1},\ldots,l_{n}\}\subseteq\texttt{Lit}^{\uparrow}$, then we have $\texttt{Lit}^{\uparrow}=\texttt{Lit}^{\uparrow}\cup l_{n+1}$ and $\texttt{Lit}^{\downarrow}=\texttt{Lit}^{\downarrow}\cup\neg l_{n+1}$.
 Example 3.
-We next illustrate the graph construction and the corresponding FC procedure for the rule set $\Gamma_{\psi}=\{\psi_{1},\psi_{2}\}$ under the Level-2 interpretation given in Example 1.
+We now illustrate the graph construction and the corresponding FC procedure for the rule set $\Gamma_{\psi}=\{\psi_{1},\psi_{2}\}$ under the Level-2 interpretation given in Example 1.
 We first obtain all the literals as $\texttt{Lit}_{\mathcal{R}}=\{p_{0},p_{1},p_{2},p_{3},p_{4}\}$ with $\texttt{Lit}^{\uparrow}=\{p_{0},p_{2}\}$ and $\texttt{Lit}^{\downarrow}=\{p_{1}\}$, construct the literal nodes correspondingly, and construct the LHS node set as $\{p_{0}\&p_{2},p_{1}\&p_{2}\}$.
-Then, directed edges are added as follows: $p_{0}\rightarrow p_{0}\&p_{2}$, $p_{2}\rightarrow p_{0}\&p_{2}$, $p_{0}\&p_{2}\rightarrow p_{3}$, $p_{1}\rightarrow p_{1}\&p_{2}$, $p_{2}\rightarrow p_{1}\&p_{2}$, $p_{1}\&p_{2}\rightarrow p_{4}$, resulting an intial graph.
+Then, directed edges are added as follows: $p_{0}\rightarrow p_{0}\&p_{2}$, $p_{2}\rightarrow p_{0}\&p_{2}$, $p_{0}\&p_{2}\rightarrow p_{3}$, $p_{1}\rightarrow p_{1}\&p_{2}$, $p_{2}\rightarrow p_{1}\&p_{2}$, $p_{1}\&p_{2}\rightarrow p_{4}$, resulting an initial graph.
 Next, we check the logic consistency.
-For the LHS node $p_{0}\&p_{2}$, since $\{p_{0},p_{2}\}\subseteq\texttt{Lit}^{\uparrow}$, we can update $\texttt{Lit}^{\uparrow}$ as $\texttt{Lit}^{\uparrow}=\texttt{Lit}^{\uparrow}\cup p_{3}=\{p_{0},p_{2},p_{3}\}$ and $\texttt{Lit}^{\downarrow}=\texttt{Lit}^{\downarrow}\cup\neg p_{3}=\{p_{1},\neg p_{3}\}$, where “$p_{3}={\sf True}$” is the new knowledge derived.
-For the LHS node $p_{1}\&p_{2}$, since $p_{1}\notin\texttt{Lit}^{\uparrow}$, there is no update happens to $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\uparrow}$; consequently, no new knowledge is inferred.
+For the LHS node $p_{0}\&p_{2}$, since $\{p_{0},p_{2}\}\subseteq\texttt{Lit}^{\uparrow}$, we update $\texttt{Lit}^{\uparrow}=\texttt{Lit}^{\uparrow}\cup p_{3}=\{p_{0},p_{2},p_{3}\}$ and $\texttt{Lit}^{\downarrow}=\texttt{Lit}^{\downarrow}\cup\neg p_{3}=\{p_{1},\neg p_{3}\}$, where $p_{3}={\sf True}$” is the new knowledge derived.
+For the LHS node $p_{1}\&p_{2}$, since $p_{1}\notin\texttt{Lit}^{\uparrow}$, there is no update on $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\uparrow}$, hence no new knowledge inferred.
 3.4 Query generation
@@
 Since the target LLM returns $\sf False$}, an inconsistency is detected, indicating a diagnostic error in the target LLM.
+Clarification.
+The soundness of this work—the ability to detect inconsistencies with respect to domain-specific constraints—depends on the accurate and faithful encoding of domain knowledge by the expert.
+Thus, the validity of RvLLM hinges on the alignment between the specification given in ESL and the underlying domain knowledge, and any misrepresentations in this encoding may compromise the soundness of the method.
 4 Experiments
-To evaluate the effectiveness of RvLLM, we apply it to the runtime verification of LLMs across three representative tasks that require domain-specific knowledge: violation detection against Singapore Rapid Transit Systems Act [1], numerical comparison, and inequality solving.
+To evaluate the effectiveness of RvLLM, we apply it to the runtime verification of LLMs across three representative tasks that require domain-specific knowledge: violation detection against Singapore Rapid Transit Systems Act [SGMRT], numerical comparison, and inequality solving.
 For a comprehensive evaluation, we include both SOTA and non-SOTA LLMs as our benchmark, including Qwen 2.5 (max, plus, turbo, 7B, 14B, 32B, 72B), GPT (4.1, 4.1 mini, 4.1 nano), Gemini 2.0 (Flash, Flash Lite), and DeepSeek-V3.
@@
 A dedicated guideline for designing interpretation abstraction prompts for the perception agent is provided in the appendices, along with additional experimental details and descriptions of the datasets and specifications used throughout the experiments.
+To ensure a comprehensive evaluation, we also compare RvLLM with other runtime approaches, including SelfCheckGPT [manakul2023selfcheckgpt_factual] and a prompt-based augmentation method that explicitly encodes domain-specific constraints within prompts.
+Due to the computational overhead and API costs associated with large-scale evaluations, we restrict our comparison to cost-efficient model variants.
+Additional experimental results and missing implementation details/justifications can be found in [zhang2025rvllm].
 4.1 Case study 1: violation detection against Singapore Rapid Transit Systems Act
 For this case study, we apply RvLLM under the Level-1 interpretation to evaluate the LLM’s responses in detecting violations within contextual scenarios against the Singapore Rapid Transit Systems Act.
-In this capacity, RvLLM can be regarded as a complementary mechanism to enhance the LLM’s ability in detecting the law violations in this study.
-We conduct experiments using contextual scenarios derived from [53], consisting of 304 cases in total–280 labeled as involving violations (unsafe) and 24 as not (safe).
+In this capacity, RvLLM can be regarded as a complementary mechanism to enhance the LLM’s capability in detecting the law violations in this study.
+We conduct experiments using contextual scenarios derived from [wang2024ali], consisting of 304 cases in total–281 labeled as involving violations (unsafe) and 23 as not (safe).
 To incorporate relevant domain knowledge, we carefully encode 31 of the 53 regulations from Singapore Rapid Transit System Act into ESL specifications.
@@
 We also find that, among all evaluated models, DeepSeek-V3 achieves the highest TPR in both standalone and integrated settings, demonstrating superior detection capabilities in this domain-specific task and good performance on language understanding.
-Table 1: Violation detection results of LLMs with and without RvLLM against the Singapore Rapid Transit Systems Act [1].
+Figure 3 gives the comparison results of RvLLM against other baseline methods.
+We observe that SelfCheckGPT often yields the highest TPR gains, as it labels nearly all outputs as “hallucinations”, resulting in near-zero TNR.
+This outcome is expected since SelfCheckGPT relies solely on output stability without leveraging domain knowledge, making it more suitable for open-domain generation.
+The prompt augmentation method (also known as LLM-as-a-judge) achieves TPR performance comparable to RvLLM but accompanied by a non-negligible TNR decrease.
+In contrast, RvLLM achieves a better balance between TPR and TNR, yielding more reliable and interpretable results.
+Table 1: Violation detection results of LLMs with and without RvLLM against the Singapore Rapid Transit Systems Act [SGMRT]. All percentage values are reported rounded to one decimal place.
 LLM LLM + RvLLM Performance of RvLLM
 Target LLMs TPR TNR TPR TNR TPR TNR Time (s)
-Qwen max [56] 56.2% 91.3% 86.1% 91.3% 29.9% $\uparrow$ 0 3.00
-Qwen plus [56] 58.4% 1 81.9% 1 23.5% $\uparrow$ 0 5.32
-Qwen turbo [56] 39.3% 1 75% 1 35.7% $\uparrow$ 0 1.64
-Qwen 2.5 (7B) [56] 16.0% 1 61.2% 87.0% 45.2% $\uparrow$ 13% $\downarrow$ 2.28
-Qwen 2.5 (14B) [56] 38.8% 95.7% 80.4% 87.0% 41.6% $\uparrow$ 8.7% $\downarrow$ 3.11
-Qwen 2.5 (32B) [56] 56.6% 95.7% 87.5% 95.7% 30.9% $\uparrow$ 0 2.55
-Qwen 2.5 (72B) [56] 57.3% 1 80.4% 95.7% 23.1% $\uparrow$ 4.3% $\downarrow$ 2.57
-GPT 4.1 [3] 57.7% 95.7% 81.1% 91.3% 23.4% $\uparrow$ 4.3% $\downarrow$ 3.11
-GPT 4.1 mini [3] 39.1% 95.7% 65.1% 95.7% 26% $\uparrow$ 0 3.94
-GPT 4.1 nano [3] 11.8% 1 30.0% 1 18.2% $\uparrow$ 0 1.82
-Gemini 2.0 Flash [21] 37.0% 1 87.2% 82.6% 50.2% $\uparrow$ 17.4% $\downarrow$ 2.69
-Gemini 2.0 Flash Lite [21] 55.1% 95.7% 79.7% 95.7% 24.6% $\uparrow$ 0 1.72
-DeepSeek-V3 [39] 78.3% 87.0% 94.0% 82.6% 15.7% $\uparrow$ 4.4% $\downarrow$ 15.84
+Qwen max [yang2024qwen2] 56.2% 91.3% 86.1% 91.3% +29.9% 0 3.00
+Qwen plus [yang2024qwen2] 58.4% 1 81.9% 1 +23.5% 0 5.32
+Qwen turbo [yang2024qwen2] 39.1% 1 74.7% 1 +35.6% 0 1.64
+Qwen 2.5 (7B) [yang2024qwen2] 16.0% 1 61.2% 87.0% +45.2% -13.0% 2.28
+Qwen 2.5 (14B) [yang2024qwen2] 38.8% 95.7% 80.4% 87.0% +41.6% -8.7% 3.11
+Qwen 2.5 (32B) [yang2024qwen2] 56.6% 95.7% 87.5% 95.7% +31.0% 0 2.55
+Qwen 2.5 (72B) [yang2024qwen2] 57.3% 1 80.4% 95.7% +23.1% -4.3% 2.57
+GPT 4.1 [achiam2023gpt] 57.7% 95.7% 81.1% 91.3% +23.5% -4.3% 3.11
+GPT 4.1 mini [achiam2023gpt] 39.1% 95.7% 65.1% 95.7% +26.0% 0 3.94
+GPT 4.1 nano [achiam2023gpt] 11.7% 1 29.9% 1 +18.1% 0 1.82
+Gemini 2.0 Flash [geminiteam2024gemini] 37.0% 1 87.2% 82.6% +50.2% -17.4% 2.69
+Gemini 2.0 Flash Lite [geminiteam2024gemini] 54.8% 95.7% 79.0% 95.7% +24.2% 0 1.72
+DeepSeek-V3 [liu2024deepseek] 78.3% 87.0% 94.0% 82.6% +15.7% -4.3% 15.84
+(a) The increase of TPR across models.
+(b) The decrease of TNR across models.
+Figure 3: Performance comparison of different methods in the violation detection task, where QW, GF/GF-Lite, and DS-V3 denote Qwen, Gemini 2.0 Flash/Flash Lite, and DeepSeek-V3, respectively.
 4.2 Case study 2: numerical comparison problems
@@
 Again, we use the same model to serve both the target LLM and the perception agent.
-Table 2: Runtime verification results of various LLMs on numerical comparison tasks.
-Target LLMs Con. Incon. Fail Time (s)
-Qwen max [56] 93 7 0 5.03
-Qwen plus [56] 99 1 0 7.01
-Qwen turbo [56] 94 6(2) 0 2.93
-Qwen 2.5 (7B) [56] 0 0 100 N/A
-Qwen 2.5 (14B) [56] 90 10(7) 0 5.09
-Qwen 2.5 (32B) [56] 98 2 0 4.98
-Qwen 2.5 (72B) [56] 96 4 0 6.31
-GPT 4.1 [3] 98 2 0 4.84
-GPT 4.1 mini [3] 92(1) 8 0 5.37
-GPT 4.1 nano [3] 90(3) 10 0 3.73
-Gemini 2.0 Flash [21] 93 7 0 4.94
-Gemini 2.0 Flash Lite [21] 52 44 4 4.35
-DeepSeek-V3 [39] 69 31(5) 0 30.87
 Results.
@@
 Among all tested models, Gemini 2.0 Flash Lite and DeepSeek-V3 show the weakest diagnostic performance, and Qwen 2.5 (7B) has the worst language understanding capabilities, failing all interpretation abstraction.
-Given its strong performance in the previous violation detection task, we conjecture that DeepSeek-V3 is better suited for language understanding tasks rather than those requiring rigorous logical reasoning.
+Table 3 presents the performance comparison across methods.
+Consistent with the findings in case study 1, we find that RvLLM achieves the best balance, maintaining both high true positive and high true negative rates among all evaluated approaches.
+Table 2: Verification results by RvLLM over various LLMs for numerical comparison tasks.
+Target LLMs Con. Incon. Fail Time (s)
+Qwen max [yang2024qwen2] 93 7 0 5.03
+Qwen plus [yang2024qwen2] 99 1 0 7.01
+Qwen turbo [yang2024qwen2] 94 6(2) 0 2.93
+Qwen 2.5 (7B) [yang2024qwen2] 0 0 100 N/A
+Qwen 2.5 (14B) [yang2024qwen2] 90 10(7) 0 5.09
+Qwen 2.5 (32B) [yang2024qwen2] 98 2 0 4.98
+Qwen 2.5 (72B) [yang2024qwen2] 96 4 0 6.31
+GPT 4.1 [achiam2023gpt] 98 2 0 4.84
+GPT 4.1 mini [achiam2023gpt] 92(1) 8 0 5.37
+GPT 4.1 nano [achiam2023gpt] 90(3) 10 0 3.73
+Gemini 2.0 Flash [geminiteam2024gemini] 93 7 0 4.94
+Gemini 2.0 Flash Lite [geminiteam2024gemini] 52 44 4 4.35
+DeepSeek-V3 [liu2024deepseek] 69 31(5) 0 30.87
+Table 3: Performance comparison of different methods for numerical comparison tasks.
+Target LLMs Methods TPR TNR
+SelfCheckGPT 0 96.7%
+Qwen plus [yang2024qwen2] LLM-as-a-judge 100% 0
+RvLLM 100% 100%
+SelfCheckGPT 100% 0
+GPT 4.1 mini [achiam2023gpt] LLM-as-a-judge 0 100%
+RvLLM 88.9% 100%
+SelfCheckGPT 87.5% 0
+Gemini 2.0 Flash [geminiteam2024gemini] LLM-as-a-judge 0 100%
+RvLLM 100% 100%
 4.3 Case study 3: inequality solving problems
-For this case study, we employ RvLLM under Level-1 interpretation to verify the step-by-step reasoning process of LLMs in inequality solving tasks.
-We compile a dataset of 40 inequality questions sourced from A-Level H2 Mathematics examination papers [2] and carefully design three specifications to serve as domain knowledge that RvLLM uses for verification.
+For this case study, we employ RvLLM under Level-1 interpretation to verify the step-by-step reasoning process of LLMs in inequality-solving tasks.
+We compile a dataset of 40 inequality questions sourced from A-Level H2 Mathematics examination papers [TestpaperJC] and carefully design three specifications to serve as domain knowledge that RvLLM uses for verification.
 These specifications consist of basic algebraic inequality properties taught at the junior college level, targeting three common LLM reasoning errors in inequality solving: incorrect factorization, flawed interval analysis, and omission of endpoint or critical point checks.
@@
 Column GT gives the ground truth number of incorrect responses attributable to the rule specifications.
-It is important to note that we define only three specifications as domain knowledge for inequality-solving tasks, which is insufficient to comprehensively cover all reasoning patterns required for such problems.
-As a result, it is expected that RvLLM fails to detect a portion of the errors due to the incompleteness of the specifications.
+It is important to note that we define only three specifications as domain knowledge for inequality-solving tasks, which is insufficient to cover all reasoning patterns required for such problems comprehensively.
+As a result, it is expected that RvLLM fails to detect a portion of the errors due to the lack of knowledge.
 The final two columns give the true positive rate of RvLLM with three specifications.
-We find that RvLLM achieves a TPR of up to 52.2% when using DeepSeek-V3 as the perception agent.
-This result highlights not only the potential of RvLLM in detecting reasoning errors (even with incomplete domain knowledge), but also underscores the importance of employing a powerful perception agent to enhance the overall performance.
-Note that, although the TPR remains below 50% for most of the models, it can reach significantly higher values when considering only the ground truth cases listed in Column GT, indicating the potential for performance improvement if more domain knowledge with respect to the inequality solving is incorporated.
+We find that RvLLM achieves a TPR of up to 50% when using DeepSeek-V3 as the perception agent.
+This result underscores both RvLLM’s potential to detect reasoning errors—even with limited domain knowledge—and the critical role of a robust perception agent in boosting overall performance.
+Although TPR remains below 50% for most models, they rise markedly when restricted to ground truth cases (Column GT), suggesting substantial gains are possible with richer domain knowledge in inequality solving.
 Table 4: Verification results by RvLLM in inequality solving tasks with three specifications, defining rules related to factorization, interval analysis, and endpoints/critical points checking.
@@
 Target LLMs Incorrect DS-V3 QW-32B GT DS-V3 QW-32B GT DS-V3 QW-32B GT DS-V3 QW-32B
-Qwen max [56] 12 1 1 1 1 0 1 2 2 3 33.3% 25%
-Qwen plus [56] 10 1 1 2 0 0 0 0 1 2 10% 20%
-Qwen turbo [56] 22 0 0 1 1 0 1 3 2 4 18.2% 9.1%
-Qwen 2.5 (7B) [56] 23 2 1 3 7 2 7 3 2 5 52.2% 21.7%
-Qwen 2.5 (14B) [56] 18 0 0 1 3 2(1) 3 3(1) 2 3 27.8% 16.7%
-Qwen 2.5 (32B) [56] 16 1 0 1 3 2 4 4 2 5 50% 25%
-Qwen 2.5 (72B) [56] 15 1 0 1 4 3 4 0 0 1 33.3% 20
-GPT 4.1 [3] 5 1 1(1) 1 0 0 0 0 0 0 20% 0
-GPT 4.1 mini [3] 2 0 0 0 0 0 0 1 0 1 50% 0
-GPT 4.1 nano [3] 10 0 0 0 1 0 1 1 0 1 20% 0
-Gemini 2.0 Flash [21] 5 0 0 0 1 0 1 0 0 0 20% 0
-Gemini 2.0 Flash Lite [21] 6 1 1 2 0 0 0 0 0 1 16.7% 16.7%
+Qwen max [yang2024qwen2] 12 1 1 1 1 0 1 2 2 3 33.3% 25%
+Qwen plus [yang2024qwen2] 10 1 1 2 0 0 0 0 1 2 10% 20%
+Qwen turbo [yang2024qwen2] 22 0 0 1 1 0 1 3 2 4 18.2% 9.1%
+Qwen 2.5 (7B) [yang2024qwen2] 25 2 1 3 7 2 7 3 2 5 48% 20%
+Qwen 2.5 (14B) [yang2024qwen2] 18 0 0 1 3 2(1) 3 3(1) 2 3 27.8% 16.7%
+Qwen 2.5 (32B) [yang2024qwen2] 16 1 0 1 3 2 4 4 2 5 50% 25%
+Qwen 2.5 (72B) [yang2024qwen2] 15 1 0 1 4 3 4 0 0 1 33.3% 20
+GPT 4.1 [achiam2023gpt] 5 1 1(1) 1 0 0 0 0 0 0 20% 0
+GPT 4.1 mini [achiam2023gpt] 2 0 0 0 0 0 0 1 0 1 50% 0
+GPT 4.1 nano [achiam2023gpt] 10 0 0 0 1 0 1 1 0 1 20% 0
+Gemini 2.0 Flash [geminiteam2024gemini] 5 0 0 0 1 0 1 0 0 0 20% 0
+Gemini 2.0 Flash Lite [geminiteam2024gemini] 6 1 1 2 0 0 0 0 0 1 16.7% 16.7%
+5 Related work
+LLM testing.
+Research in this area aims to develop comprehensive benchmarks and systematic evaluation frameworks to assess LLM across general attributes such as robustness, fairness, and factual consistency [liu2023trustworthy, liang2022holistic, wang2023decodingtrust].
+Robustness studies examine model resilience to adversarial input, paraphrasing, or noise [cai2024benchlmm, tu2023many, zhang2024multitrust], while fairness evaluations assess bias with respect to sensitive attributes like gender or race [li2024red, zhang2024multitrust].
+Factual consistency detection, or hallucination detection, focuses on identifying instances in which LLMs produce plausible yet factually incorrect or unsupported claims [lin2022truthfulqa, li2023halueval, li2024drowzee].
+Although large-scale benchmarks like HELM [liang2022holistic] and BIG-bench [srivastava2022beyond] offer task suites and metrics for performance evaluation, a critical challenge remains—the dynamic and open-ended nature of real-world deployments complicates the construction of exhaustive and representative test suites.
+Runtime verification of LLMs.
+Runtime verification [bartocci2018introduction, falcone2013tutorial] is a dynamic analysis technique that ensures certified system behavior during execution by monitoring traces against formal specifications.
+Owing to its lightweight nature, runtime verification has inspired multiple efforts to verify LLM-generated outputs at runtime [cohen2023lm_factual, besta2024checkembed_factual, manakul2023selfcheckgpt_factual, chen2024complex_factual, laban2022summac_faithful, luo2023chatgpt_faithful, laban2023llms_faithful].
+However, existing approaches center on open-domain settings, where specifications are often ambiguous and informal.
+For instance, SelfCheckGPT [manakul2023selfcheckgpt_factual] and CHECKEMBED [besta2024checkembed_factual] operationalize specifications in terms of output stability.
+More recently, [cheng2024formal] proposes a fairness monitoring approach leveraging precisely defined properties in formal specifications.
+These specifications, informed by linear temporal logic [pnueli1977temporal] and its bounded metric variant [alur1996benefits], signal a shift towards formal methods for more accurate and dependable runtime verification of LLM behaviors.
+However, these works primarily focus on general properties and are not well-suited to domain-specific constraints for specialized tasks.
+Expert systems with NLP.
+Rule-based expert systems [daniel1997cadiag, denning1986towards, durkin1996expert] have effectively tackled domain-specific problems by coupling knowledge bases with inference engines to emulate expert reasoning for decades of years.
+To enhance usability, in the 70s, researchers integrated natural language processing (NLP), enabling interaction via simplified language, e.g., SHRDLU [winograd1971procedures] in virtual environments and MedLEE [krauthammer2001knowledge] in clinical text analysis.
+These systems relied on symbolic parsing and controlled vocabularies but suffered from fragile parsing, limited lexical coverage, and ambiguity, often requiring constrained input or extensive user guidance.
+Nevertheless, integrating NLP with rule-based reasoning provides a key proof-of-concept for interactive, user-friendly AI, shaping later knowledge-based systems and natural language interfaces.
+Recently, the rise of LLMs has reinvigorated this paradigm.
+In this work, we employ a perception LLM to translate the target LLM’s context and outputs into formal representations for backend reasoning.
+An improved version could explore other extensions of Horn clauses beyond those considered here to enhance expressiveness and completeness [babka2013complexity].
 6 Conclusion
@@
 For each task, we develop tailored specifications grounded in relevant domain expertise.
-Experimental results indicate that RvLLM effectively leverages expert-defined domain knowledge to enable precise runtime verification of LLMs.
-However, despite its good performance, the current framework suffers from limited expressiveness, restricting its ability to encode more complex domain knowledge and behavioral constraints for LLMs.
-Furthermore, the reasoning mechanism remains limited to a single strategy, as embodied in the forward chaining module.
-Future work will focus on enriching the specification language with additional operators to enhance its expressiveness and investigating the integration of more reasoning strategies.
-This is intended to enable a more nuanced and comprehensive inference process, thereby enhancing overall verification capabilities.
+Experimental results indicate that RvLLM effectively leverages expert-defined domain knowledge to enable precise runtime verification of LLMs and strike a good balance between TPR and TNR, compared to existing runtime methods.
+However, despite its good performance, the current framework suffers from limited expressiveness by ESL, restricting its ability to encode more complex domain knowledge and behavioral constraints for LLMs.
+Future work will focus on enriching the specification language with additional operators to improve its expressiveness and generalization.
+7 Acknowledgments
+This work was partially funded by the Ministry of Education, Singapore, under its Academic Research Fund Tier 3 (MOET32020-0003).
+Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
+Appendix A Further Examples of the FC Algorithm
+In this section, we illustrate the graph construction and the corresponding forward chaining process by a more complicated example. Consider the following rule set $\Gamma_{\mathcal{R}}$:
+$\{\neg A\wedge B\Rightarrow\neg M,A\wedge\neg B\Rightarrow\neg N,A\wedge\neg B\Rightarrow P,\neg M\wedge N\Rightarrow Q,\neg N\wedge P\Rightarrow K\}$
+We first obtain all the literals as $\texttt{Lit}_{\mathcal{R}}=\{A$, $\neg A$, $B$, $\neg B$, $\neg M$, $N$, $\neg N$, $P$, $Q$, K}K\}, construct the literal nodes correspondingly, and construct the LHS nodes as $\{\neg A\wedge B,A\wedge\neg B,\neg M\wedge N,\neg N\wedge P\}$ according to $\Gamma_{\mathcal{R}}$.
+Then, directed edges are added to represent the relationships between LHS and RHS, leading to the construction of the graph illustrated in Figure 4(a).
+Assume that the initially defined literals are $\texttt{Lit}=\{A,B,\neg M\}$ with truth value as $A=\neg M={\tt True}$ and $B={\tt False}$, then we have the intial node set with value $\sf True$} as $\texttt{Lit}^{\uparrow}=\{A,\neg B,\neg M\}$ (marked in green) and the initial false node set with value $\sf False$} as $\texttt{Lit}^{\downarrow}=\{\neg A,B\}$ (marked in red) in Figure 4(b).
+Since $\{A,\neg B\}\subseteq\texttt{Lit}^{\uparrow}$, by applying one time of modus ponens inference rule on $A\wedge\neg B\Rightarrow\neg N$ and $A\wedge\neg B\Rightarrow P$, we can update $\texttt{Lit}^{\uparrow}$ as $\{A,\neg B,\neg M,\neg N,P\}$ and $\texttt{Lit}^{\downarrow}$ as $\{\neg A,B,N\}$, as shown in Figure 4(c).
+Similarly, based on $\{\neg N,P\}\subseteq\texttt{Lit}^{\uparrow}$, we update the nodes set $\texttt{Lit}^{\uparrow}$ as $\texttt{Lit}^{\uparrow}\cup\{K\}$ and remain $\texttt{Lit}^{\downarrow}$ unchanged, as shown in Figure 4(d).
+Finally, with $\{\neg M,K\}\subseteq\texttt{Lit}^{\uparrow}$ for the LHS node $\neg M\wedge K$, we update the node set $\texttt{Lit}^{\uparrow}$ as $\texttt{Lit}^{\uparrow}\cup\{N\}=\{A,\neg B,\neg M,\neg N,P,K,N\}$.
+Since $\texttt{Lit}^{\uparrow}\cup\texttt{Lit}^{\downarrow}=\{N\}$, an inconsistency is identified.
+Note that if $\neg M$ is removed from the initially defined literal set Lit, then there will be no inconsistency, and we will just obtain values for the undefined literals $\{\neg N,N,P,K\}$ following this FC procedure.
+(a) Initial graph $G$ construct following the three steps.
+(b) $G$ with initially assigned literals and the corresponding nodes with values $\sf True$} or $\sf False$}.
+(c) $G$ after one time of modus ponens.
+(d) $G$ after two times of modus ponens.
+Figure 4: The initial graph $G$ is constructed in (4(a)).
+The literal node sets $\texttt{Lit}^{\uparrow}$ and $\texttt{Lit}^{\downarrow}$ are updated step by step (from (4(b)) to (4(d))), following the modus ponens inference rules, and an inconsistency is detected on the literal node NN in (4(d)).
+Appendix B Dataset and experimental design in RvLLM
+To evaluate the effectiveness of RvLLM, we consider three case studies: i) violation detection under the Singapore Rapid Transit Systems Act [SGMRT], ii) numerical comparison tasks, and iii) inequality solving problems.
+In the following, we present an in-depth overview of each case study and the corresponding datasets used.
+Dataset 1.
+Singapore Rapid Transit Systems Act.
+This dataset comprises a collection of regulations governing passenger conduct on the railway.
+Among the full set of regulations, we identify and select 31 types that pertain specifically to passenger behaviors.
+These include: Trespass, Smoking, Luggage, Foldable bicycle, Bringing animal, Dangerous goods, Offensive matter, Compliance with instructions, Overcrowding, Consumption, Spitting, Causing nuisance, Loitering, Throwing missiles, Wrongfully entering, Hawking, Meddling with plant, Misuse of an escalator, Misuse of an emergency safety device, Interference with doors, Improperly removing, Causing obstruction and danger, Transferring articles, Damaging a ticket, Hand in tickets, Failure to pay, Leaving motor vehicles, Failure of a vehicle driver, Dangerous driving, Driving vehicles on certain parts, Lost property.
+For each of these regulation types, the work [wang2024ali] provides a set of scenarios–both compliant and non-compliant–that serve as our benchmark for conducting violation detection.
+In the experimental design, we first query the target LLM to determine whether a given scenario is “Safe”–that is, free of any violations of the Singapore Rapid Transit Systems Act.
+This provides a baseline for violation detection using only the LLM.
+Subsequently, we employ RvLLM, which takes the scenario solely as the domain of discourse based on a Level-1 interpretation.
+The violation detection results are then collected from RvLLM.
+For scenarios that contain actual violations, if the target LLM classifies the scenario as “Safe” (i.e., fails to detect the violation), while RvLLM identifies a violation and returns “Inconsistent”, we consider this as evidence that RvLLM improves the detection accuracy of the target LLM on that particular example.
+Dataset 2.
+Numerical comparison.
+For this dataset, we randomly generate 100 numerical comparison problems to serve as our benchmark.
+This design is motivated by well-documented errors in LLMs when handling numerical comparisons, particularly in cases where the integral parts are identical but the fractional parts differ.
+A notorious example is the comparison between $9.11$ and $9.9$, which has been widely observed to cause incorrect behavior in LLMs [compError].
+In this case, models often erroneously compare the fractional part $0.11$ with $0.9$.
+Building on this insight, we construct 100 pairs of decimal numbers that share the same integer part but differ in their fractional components, specifically in the form $x.y$ vs.
+$x.z$, where $y<10<z$.
+Example pairs include “7.47 vs.
+7.9” and “33.3 vs.
+33.27”.
+These are designed to systematically evaluate how well RvLLM and baseline methods detect inconsistency in LLMs with respect to the comparison rule.
+In the experimental design, each numerical comparison question is posed to the target LLM with the instruction: “Please answer which one is greater, but nothing else.” The LLM’s response is then collected.
+Subsequently, RvLLM treats both the original comparison question and the LLM output as the domain of discourse, and performs verification against the specification file under the Level-2 interpretation.
+Dataset 3.
+Inequality solving.
+For this case study, we collect 40 inequality-solving problems from A-Level H2 Mathematics examination papers [TestpaperJC].
+All of these problems involve a single free variable and are expressed as polynomial inequalities of degree two or higher, including examples such as fourth-degree polynomials.
+Here we give some example question used in the experiment: $\frac{3x-4x^{2}}{2x+1}\geq 0$, $\frac{x+5}{(x-3)^{2}}\geq\frac{x-7}{x(x-3)}$, and $\frac{4x^{2}-8x+2}{x-2}>\frac{1}{x^{2}}$.
+In the experimental design, each inequality-solving question is posed to the target LLM with the instruction: "Please solve the following inequality problem." The resulting step-by-step reasoning generated by the LLM is then provided as the domain of discourse for RvLLM.
+During this verification, RvLLM checks the logical consistency of the LLM’s reasoning steps against three specification files encoding fundamental algebraic properties of inequalities taught at the junior college level.
+These specifications are specifically designed to detect three common categories of reasoning errors in LLM-generated solutions: incorrect factorization, flawed interval analysis, and the omission of endpoint or critical point evaluations.
+Appendix C Prompt design in RvLLM
+In this section, we present the prompt used across all experiments.
+It is worth noting that, for more effective perception, it is recommended to design distinct prompt templates tailored to specific downstream, domain-specific tasks.
+C.1 Prompts in case study 1
+Original query. The prompt shown in Figure 5 serves as the original query template provided to the target LLM, with the testing scenario serving as the context part [[Context]].
+Task Description: (1) Please evaluate the severity for the following context: True (There is no misconduct behavior in the context according to the railway laws in Singapore) or False (There exists some misconduct in the context according to the railway laws in Singapore).
+(2) Please only output the single word ’True’ or ’False’ but nothing else.
+Context: [[Context]]
+Figure 5: The query prompt for case study 1, where [[Context]] is replaced with the corresponding testing scenario.
+Prompt for Level-1 interpretation abstraction.
+The prompt template provided in Figure 6 is designed for the perception agent to perform the Level-1 interpretation, that is, to abstract all feasible facts, or instantiations of the given predicates, based on the specification file.
+Given a specific specification file, the corresponding interpretation prompt can be generated by replacing the placeholders: i) [[Variables]] and [[Predicates]] with the ones from the specification file, and ii) [[Context]] with the relevant testing scenario.
+Task Description: You are a perception agent with NO reasoning abilities.
+(1) Get the “Objects” and “Interpretation” only based on the “Context” within the "Domain" following the way used in the "Template".
+(2) There are three values of each predicate: True, False, and Unknown.
+(3) Set “Unknown” to the predicate when there is no “explicit evidence”.
+(4) The **arguments** in each interpretation must be **from Objects**.
+(5) Only output two lists but nothing else: “Objects” and “Interpretations”.
+– – – – – Template Start – – – – – Variables: x, y Predicates: inMRT(x) := x is on MRT.
+consuming(x) := x is consuming something.
+Context: Alice and Bob are on the MRT.
+Alice is eating bread.
+Objects: {Alice, Bob, bread} Interpretations: {inMRT(Alice) = True; inMRT(Bob) = True; consuming(Alice, bread) = True; consuming(Bob, bread) = Unknown;} – – – – – Template End – – – – – – – – – – Domain Start – – – – – Variables: [[Variables]] Predicates: [[Predicates]] Context: [[Context]] – – – – – Domain End – – – – –
+Figure 6: The Level-1 interpretation abstraction prompt for case study 1.
+[[Variables]] and [[Predicates]] placeholders are substituted with the ones from the specification file.
+[[Context]] is replaced with the testing scenario.
+Prompt for new queries.
+The prompt template provided in Figure 7 is designed to issue a new query to the target LLM to question the inferred knowledge.
+In this setting, the new query must refer back to its original context, namely the testing scenario.
+Task Description: (1) Given the Context and Predicates, please check the Boolean value of the Proposition.
+(2) Return True or False only.
+(3) Only output True or False.
+Context: [[Context]] Predicates: [[Predicates]] Proposition: [[Proposition]]
+Figure 7: The query prompt for the **inferred** knowledge used in case study 1.
+[[Context]] is replaced with the testing scenario, [[Predicates]] is substituted with those from the specification file, and [[Proposition]] is replaced with the inferred knowledge expressed as instantiated predicates.
+C.2 Prompt in case study 2
+Original query. The prompt shown in Figure 8 serves as the original query template provided to the target LLM.
+Task Description: [[comp_a]] and [[comp_b]], which one is greater? Please output the result directly with nothing else.
+Figure 8: The query prompt for case study 2, where [[comp_a]] and [[comp_b]] are replaced with the specific numbers to be compared.
+Prompt for Level-1 interpretation abstraction.
+The prompt template shown in Figure 9 is designed for the perception agent to perform the interpretation abstraction for case study 2.
+The substitution of the placeholders [[Variables]] and [[Predicates]] follows the same procedure as illustrated in Figure 6.
+In this case, [[Context]] is replaced with the comparison query submitted to the target LLM and its corresponding response.
+Task Description: You are a perception agent with NO reasoning abilities.
+(1) Get the “Objects” and “Interpretation” only based on the “Context” within the "Domain" following the way used in the "Template".
+(2) There are three values of each predicate: True, False, and Unknown.
+(3) Set “Unknown” to the predicate when there is no “explicit evidence”.
+(4) The **arguments** in each interpretation must be **from Objects**.
+(5) Only output two lists but nothing else: “Objects” and “Interpretations”.
+– – – – – Template Start – – – – – Variables: x, y Predicates: IsLarger(x, y) := x is larger than y.
+Pos(x) := x is a positive number, i.e., x > 0.
+Neg(x) := x is a negative number, i.e., x < 0.
+Context: Question: 1 and 10, which one is greater?
+Answer: 10 Objects: {1, 10} Interpretations: {IsLarger(10, 1) = True; IsLarger(1, 10) = False; Pos(1) = Unknown; Neg(1) = Unknown; Pos(10) = Unknown; Neg(10) = Unknown;} – – – – – Template End – – – – – – – – – – Domain Start – – – – – Variables: [[Variables]] Predicates: [[Predicates]] Context: [[Context]] – – – – – Domain End – – – – –
+Figure 9: The Level-1 interpretation abstraction prompt for case study 2.
+[[Variables]] and [[Predicates]] are replaced with the ones from the specification file.
+[[Context]] is replaced with the comparison question together with the corresponding response generated by the target LLM.
+Prompt for Level-2 interpretation abstraction.
+The prompt template provided in Figure 10 illustrates the prompt used to generate a new instance—invoked when a Level-2 interpretation is triggered—for a predicate that was not fully instantiated.
+This process addresses the partially interpreted rule that arises at the Level-1 interpretation.
+Note that we reuse the template part from case study 1 directly, as this helps avoid unintentionally generating a fixed instance.
+[[Variables]] and [[Predicates]] are replaced with those from the specification file.
+[[Values]] is filled with the instantiated predicates and their assigned truth values, guiding the perception agent in generating the desired instances.
+Task Description: (1) Generate an instance for each variable in the predicate and satisfy the predicate value.
+You can follow the way used in “Template”.
+(2) The predicate values should be satisfied at the same time.
+(3) Only output the variable instance value with nothing else.
+– – – – – Template Start – – – – – Variables: x, y Predicates: InRailwayPremises(x, y) := a person x is in some railway premises y.
+Values: {InRailwayPremises(x, y) = True} Instance: {x = Alice; y = Railway station} – – – – – Template End – – – – – – – – – – Domain Start – – – – – Variables: [[Variables]] Predicates: [[Predicates]] Values: [[Values]] – – – – – Domain End – – – – –
+Figure 10: The Prompt for new-instance generation during Level-2 interpretation abstraction in case study 2.
+[[Variables]] and [[Predicates]] are replaced with the ones from the specification file.
+[[Values]] is replaced with the predicate along with its assigned truth value.
+Prompt for new queries.
+The prompt template provided in Figure 11 is designed to issue a new query to the target LLM to question the inferred knowledge.
+Unlike the prompt in Figure 7, the original context is unnecessary in this setting because the inferred knowledge consists solely of concrete numerical comparison results.
+Instead, the relevant predicate definitions ([[Predicates]]) and the inferred propositional knowledge ([[Propositions]]) are required.
+Task Description: You are a perception agent with NO reasoning abilities.
+(1) Given the Predicates, please check the Boolean value of the Proposition.
+(2) Return True or False only.
+(3) Only output True or False.
+Predicates: [[Predicates]] Proposition: [[Proposition]]
+Figure 11: The query prompt for the **inferred** knowledge used in case study 2.
+[[Predicates]] is replaced with the ones from the specification file.
+[[Proposition]] is replaced with the inferred knowledge, which is represented as instantiated predicates.
+C.3 Prompt in case study 3
+Original query. The prompt shown in Figure 12 serves as the original query template provided to the target LLM.
+Task Description: Please solve the following inequality exactly: [[inequality]]
+Figure 12: The query prompt for case study 3, where [[inequality]] is replaced with the specific inequality problem to be solved.
+Prompt for Level-1 interpretation abstraction.
+Due to the task complexity, we keep the predicate and variable fixed, using those specified in the specification file.
+Only the [[Context]] placeholder is replaced with the query response from the target LLM, as illustrated in Figures 13.
+Task Description: You are a perception agent with NO reasoning abilities.
+(1) Get the “Objects” and “Interpretation” ONLY based on the “Context” within the “Domain” without any reasoning procedure.
+(2) There are three values of each predicate: True, False, and Unknown.
+(3) The **arguments** in each interpretation must be **from Objects**.
+(4) Only output two lists: Objects and Interpretation.
+– – – – – Template Start – – – – – Variables: [[Variables]] Predicates: [[Predicates]] Context: ### Factor the quadratic expression -x^2 - ax + b = -(x^2 + cx + d) = -(x+e)^2 Simplify (x+a)(x-b) and get (x^2 + cx - d).
+Factor y^2 - cy + d = (y-a)(y+b) Objects: {minus_x_squared_minus_2_x_minus_1, x_squared_plus_c_x_minus_d, y_squared_minus_c_y_plus_d, x_plus_e, x_plus_a, x_minus_b, y_minus_a, y_plus_b, -1, -a, b, 1, c, d, e, a, -b, -d, -c} Interpretations: {PolyFactorizeSquareNeg(minus_x_squared_minus_a_x_plus_b, x_plus_e) = True; PolyFactorizeSquarePos(x_squared_plus_c_x_plus_d, x_plus_e) = True; PolyFactorizePos(x_squared_minus_c_x_minus_d, x_plus_a, x_minus_b) = True; PolyFactorizePos(y_squared_minus_c_y_plus_d, y_minus_a, y_plus_b) = True; IsPolyTwo(minus_x_squared_minus_a_x_plus_b, -1, -a, b, x) = True;   IsPolyTwo(x_squared_plus_c_x_plus_d, 1, c, d, x) = True; IsPolyTwo(x_squared_minus_c_x_minus_d, 1, c, -d, x) = True;   IsPolyTwo(y_squared_minus_c_y_plus_d, 1, -c, d, x) = True; IsPolyOne(x_plus_e, 1, e, x) = True;   IsPolyOne(x_plus_a, 1, a, x) = True;   IsPolyOne(x_minus_b, 1, -b, x) = True; IsPolyOne(y_minus_a, 1, -a, y) = True;   IsPolyOne(y_plus_b, 1, b, y) = True;} – – – – – Template End – – – – – – – – – – Domain Start – – – – – Context: [[Context]] – – – – – Domain End – – – – –
+Figure 13: The prompt template of interpretation abstraction used in case study 3: Factorization Error.
+[[Variables]] and [[Predicates]] are substituted with the ones from the specification file.
+[[Context]] is replaced with the target LLM’s step-by-step reasoning process.
+Prompt for new queries.
+For the factorization and interval specifications, we employ the same prompt template shown in Figure 11.
+However, for the endpoints specification, the target LLM must refer to its original step-by-step reasoning process when answering inferred knowledge related to endpoint information; therefore, the corresponding prompt template is identical to the one in Figure 7.
+Appendix D An illustration example with RvLLM
+We illustrate the usage of RvLLM through a concrete example drawn from case study 1. Consider the following regulation on bringing animals and its corresponding ESL specification file:
+  • Regulation No.8 (1) No person shall bring any animal into or upon, or allow any animal under his control to remain in or on, any part of the railway premises.
+(2) A person shall be responsible for any injury, loss or damage caused to the property or staff of the Authority or its licensee, or to any other person or property by such person or by any animal or article brought by him onto the railway premises and he shall indemnify the Authority and its licensee from and against any liability to any other person resulting therefrom.
+(3) The prohibition in paragraph (1) shall not apply to – (a) a guide dog accompanying a person with a sight or hearing impairment; or (b) a dog under the effective control of any police officer on official duty or any member of an auxiliary police force in uniform or any security officer (within the meaning of the Private Security Industry Act (Cap.
+250A)) engaged by the Authority or its licensee to provide security at the railway premises.
+"Variables": {"x", "y"},
+"Predicates": {"IsAnimal(x) := x is an animal.",
+"IsGuideDog(x) := x is a guide dog accompanying a person with a sight or hearing impairment.",
+"IsOfficerDog(x) := x is a dog under the effective control of a police officer, auxiliary police officer, or security officer on duty.",
+"CauseDamageOrInjury(x) := some person x or the animal of some person x brought with causes injury, loss, or damage to persons or property.",
+"InRailwayPremises(x) := a person x or an animal x is in some railway premises or on the railway station."},
+"Rules": {"IsAnimal(x) & InRailwayPremises(x) & not IsOfficerDog(x) => IsGuideDog(x)",
+"IsAnimal(x) & InRailwayPremises(x) & not IsGuideDog(x) => IsOfficerDog(x)"}
+Now consider the following scenario which is evaluated as “Safe” by the target LLM qwen-plus:
+“A commuter on the MRT carefully transported their pet snake in a secure, locked container designed to ensure the safety of other passengers.
+The container was opaque to minimize any potential disturbance or alarm among fellow commuters.
+The owner calmly explained to curious onlookers that the snake was being moved to a new home and assured them of the safety measures taken.
+Throughout the journey, the snake remained quiet and undetectable, causing no disruption to the train’s operation or its passengers.”
+Given the input scenario, RvLLM first invokes the perception agent (also qwen-plus in this example), to extract the objects and their associated interpretations using the prompt shown in Figure 6.
+Results are summarized in Table 5.
+Table 5: The interpretation obtained by the perception agent qwen-plus.
+objects_all “commuter”, “snake”, “container”
+“IsAnimal(snake) = True”, “IsGuideDog(snake) = False”, “IsOfficerDog(snake) = False”
+interPre_all “CauseDamageOrInjury(snake) = False”, “InRailwayPremises(commuter) = True”
+“InRailwayPremises(snake) = True”
+Recall that the specification rules provided in the “Bring Animal” regulation are as follows:
+  1. 1.
+    $\texttt{IsAnimal}(x)\ \&\ \texttt{InRailPremises(x)}\ \&\ \neg\texttt{IsOfficerDog}(x)\Rightarrow\texttt{IsGuideDog}(x)$
+  2. 2.
+    $\texttt{IsAnimal}(x)\ \&\ \texttt{InRailPremises(x)}\ \&\ \neg\texttt{IsGuideDog}(x)\Rightarrow\texttt{IsOfficerDog}(x)$
+Following the interpretation process, we obtain the propositional, rule-like formulas listed in Table 6, each formulated using instantiated predicates.
+For clarity and ease of representation, we adopt the notation summarized in Table 7 to denote the propositions that appear in these interpreted formulas.
+Table 7 also provides the truth value of each proposition, determined by the perception agent’s outputs.
+If a truth value cannot be inferred from perception, it is assigned $\sf Unknown$} under the open-world assumption.
+Finally, we obtain the rephrased propositional, rule-like formulas, as presented in Table 8.
+Table 6: The interpreted propositional rule-like formula derived from the specification rules, where formulae (i)-(iii) are obtained from the interpretation of the first rule, and formulae (iv)-(vi) are derived from the second rule.
+i $\texttt{IsAnimal}(\text{commuter})$ & $\texttt{InRailPremises}(\text{commuter})$ & $\neg\texttt{IsOfficerDog}(\text{commuter})$ $\Rightarrow$ $\texttt{IsGuideDog}(\text{commuter})$
+$i$ $\texttt{IsAnimal}(\text{snake})$ & $\texttt{InRailPremises}(\text{snake})$ & $\neg\texttt{IsOfficerDog}(\text{snake})$ $\Rightarrow$ $\texttt{IsGuideDog}(\text{snake})$
+iii $\texttt{IsAnimal}(\text{container})$ & $\texttt{InRailPremises}(\text{container})$ & $\neg\texttt{IsOfficerDog}(\text{container})$ $\Rightarrow$ $\texttt{IsGuideDog}(\text{container})$
+iv $\texttt{IsAnimal}(\text{commuter})$ & $\texttt{InRailPremises}(\text{commuter})$ & $\neg\texttt{IsGuideDog}(\text{commuter})$ $\Rightarrow$ $\texttt{IsOfficerDog}(\text{commuter})$
+v $\texttt{IsAnimal}(\text{snake})$ & $\texttt{InRailPremises}(\text{snake})$ & $\neg\texttt{IsGuideDog}(\text{snake})$ $\Rightarrow$ $\texttt{IsOfficerDog}(\text{snake})$
+vi $\texttt{IsAnimal}(\text{container})$ & $\texttt{InRailPremises}(\text{container})$ & $\neg\texttt{IsGuideDog}(\text{container})$ $\Rightarrow$ $\texttt{IsOfficerDog}(\text{container})$
+Table 7: The propositional representations of the interpreted predicates, where $\top$, $\bot$, and $\sf Unk.$} denote the truth value as $\sf True$}, $\sf False$}, and $\sf Unknown$}, respectively.
+$p_{0}$: $\texttt{IsAnimal}(\text{commuter})={\sf Unk.}$ $p_{6}$: $\texttt{IsOfficerDog}(\text{commuter})={\sf Unk.}$
+$p_{1}$: $\texttt{IsAnimal}(\text{snake})=\top$ $p_{7}$: $\texttt{IsOfficerDog}(\text{snake})=\bot$
+$p_{2}$: $\texttt{IsAnimal}(\text{container})={\sf Unk.}$ $p_{8}$: $\texttt{IsOfficerDog}(\text{container})={\sf Unk.}$
+$p_{3}$: $\texttt{InRailPremises}(\text{commuter})=\top$ $p_{9}$: $\texttt{IsGuideDog}(\text{commuter})={\sf Unk.}$
+$p_{4}$: $\texttt{InRailPremises}(\text{snake})=\top$ $p_{10}$: $\texttt{IsGuideDog}(\text{snake})=\bot$
+$p_{5}$: $\texttt{InRailPremises}(\text{container})={\sf Unk.}$ $p_{11}$: $\texttt{IsGuideDog}(\text{container})={\sf Unk.}$
+Table 8: The rephrased propositional rule-like formula as well as the truth value of the propositions.
+i $p_{0}$ & $p_{3}$ & $\neg p_{6}$ $\Rightarrow$ $p_{9}$ $p_{0}=p_{6}=p_{9}={\sf Unk.}$, $p_{3}=\top$
+$i$ $p_{1}$ & $p_{4}$ & $\neg p_{7}$ $\Rightarrow$ $p_{10}$ $p_{1}=p_{4}=\top$, $p_{7}=p_{10}=\bot$
+iii $p_{2}$ & $p_{5}$ & $\neg p_{8}$ $\Rightarrow$ $p_{11}$ $p_{2}=p_{5}=p_{8}=p_{11}={\sf Unk.}$
+iv $p_{0}$ & $p_{3}$ & $\neg p_{9}$ $\Rightarrow$ $p_{6}$ $p_{0}=p_{9}=p_{6}={\sf Unk.}$, $p_{3}=\top$
+v $p_{1}$ & $p_{4}$ & $\neg p_{10}$ $\Rightarrow$ $p_{7}$ $p_{1}=p_{4}=\top$, $p_{10}=p_{7}=\bot$
+vi $p_{2}$ & $p_{5}$ & $\neg p_{11}$ $\Rightarrow$ $p_{8}$ $p_{2}=p_{5}=p_{11}=p_{8}={\sf Unk.}$
+After conducting the forward chaining procedure, we identify an inconsistency arising from the propositional rules (ii) and (v). The corresponding inconsistent knowledge is:
+$\texttt{IsAnimal}(\text{snake})\&\texttt{InRailwayPremises}(\text{snake})\&\neg\texttt{IsOfficerDog}(\text{snake})\Rightarrow\texttt{IsGuideDog}(\text{snake})$
+$\texttt{IsAnimal}(\text{snake})\&\texttt{InRailwayPremises}(\text{snake})\&\neg\texttt{IsGuideDog}(\text{snake})\Rightarrow\texttt{IsOfficerDog}(\text{snake})$
+Analysis.
+It is worth noting that the target LLM does not identify any violation, as the context appears to indicate no disruption to the train or its passengers.
+However, according to the strict regulations on bringing animals aboard, only guide dogs and official dogs are permitted on the premises.
+By leveraging domain-specific knowledge encoded in the ESL-based specification, RvLLM successfully detects the violation that the LLM overlooks.
+Appendix E A comprehensive comparative analysis of baseline methods
+In addition to evaluating the performance of RvLLM against raw LLM outputs, we further conduct comparative experiments with other established baseline methods.
+Specifically, we compare RvLLM with: i) SelfCheckGPT [manakul2023selfcheckgpt_factual] and ii) a prompt-augmentation approach, in which domain-specific natural language constraints are explicitly encoded into the prompt of another LLM acting as a judge—commonly referred to as the LLM-as-a-judge paradigm.
+Implementation details.
+For a fair comparison, we employ DeepSeek-V3 uniformly as the perception agent in RvLLM and as the judge model in the LLM-as-a-judge baseline.
+For SelfCheckGPT, we adopt the authors’ recommended settings by enabling the NLI option (by default) with a hallucination threshold of 0.001, and using a sampling count of $4$.
+In addition, due to the considerable computational overhead and API costs of large-scale evaluation, we adopt cost-efficient model variants (i.e., models with a lower-priced API) across all comparative experiments.
+The comparative results are summarized below.
+For case study 3, we additionally report the average token usage per task.
+Table 9: Violation detection performance in case study 1. Values show changes in TPR and TNR relative to raw LLM outputs under each method.
+Target LLMs Methods TPR Inc. TNR Dec.
+SelfCheckGPT +41.6% -100%
+Qwen plus [yang2024qwen2] LLM-as-a-judge +31.0% -56.5%
+RvLLM +23.5% 0
+SelfCheckGPT +60.9% -100%
+Qwen-turbo [yang2024qwen2] LLM-as-a-judge +38.1% -47.8%
+RvLLM +35.6% 0
+SelfCheckGPT +60.9% -95.7%
+GPT 4.1 mini [achiam2023gpt] LLM-as-a-judge +22.4% -17.4%
+RvLLM +26.0% 0
+SelfCheckGPT +63.0% -100%
+Gemini 2.0 Flash [geminiteam2024gemini] LLM-as-a-judge +7.8% -13.0%
+RvLLM +50.2% -17.4%
+SelfCheckGPT +45.2% -95.7%
+Gemini 2.0 Flash Lite [geminiteam2024gemini] LLM-as-a-judge +18.1% -21.7%
+RvLLM +24.2% 0
+SelfCheckGPT +21.7% -87.0%
+DeepSeek-V3 [liu2024deepseek] LLM-as-a-judge +9.6% -65.2%
+RvLLM +15.7% -4.3%
+Results.
+The comparison results for case study 1 are given in Table 9.
+We observe that RvLLM achieves more stable performance across models, consistently maintaining a relatively small TNR drop when improving TPR.
+In contrast, SelfCheckGPT often yields the best TPR gains but tends to classify nearly all outputs as “hallucinations” in our domain-specific setting, resulting in near-zero TNR.
+This behavior is expected, as SelfCheckGPT relies solely on output stability and does not incorporate domain knowledge.
+It neither encodes nor evaluates domain-specific rules, constraints, or context-sensitive reasoning, making it more appropriate for open-domain generation tasks.
+For the LLM-as-a-judge method, we find that it achieves TPR gains comparable to those of RvLLM.
+However, its performance is also unreliable, with TPR improvements typically accompanied by substantial drops in TNR.
+In contrast, RvLLM maintains a more favorable balance between TPR and TNR.
+This is because RvLLM leverages the LLM primarily for text understanding, while delegating rule-based, domain-specific reasoning to an external reasoning engine, resulting in more reliable and interpretable outcomes.
+The results of the remaining two comparative experiments are presented in Tables 10 and 11.
+Similar to the results in case study 1, we find that for both SelfCheckGPT and LLM-as-a-judge, a high TPR (resp.
+TNR) is typically accompanied by a low TNR (resp.
+TPR), as these methods tend to label all test cases as positive (resp.
+negative).
+Recall that a positive instance indicates the presence of a violation of the specification rule within the LLM’s context.
+We also note that the relatively lower TPR of RvLLM in case study 3 stems from the limited domain-specific knowledge currently encoded.
+As more relevant knowledge is incorporated, the TPR of RvLLM is expected to improve while still maintaining a high TNR.
+Table 10: Performance of different methods for numerical comparison tasks in case study 2.
+Target LLMs Methods TPR TNR
+SelfCheckGPT 0 96.7%
+Qwen plus [yang2024qwen2] LLM-as-a-judge 100% 0
+RvLLM 100% 100%
+SelfCheckGPT 100% 7.3%
+Qwen-turbo [yang2024qwen2] LLM-as-a-judge 100% 0
+RvLLM 100% 97.9%
+SelfCheckGPT 100% 0
+GPT 4.1 mini [achiam2023gpt] LLM-as-a-judge 0 100%
+RvLLM 88.9% 100%
+SelfCheckGPT 87.5% 0
+Gemini 2.0 Flash [geminiteam2024gemini] LLM-as-a-judge 0 100%
+RvLLM 100% 100%
+SelfCheckGPT 100% 0
+Gemini 2.0 Flash Lite [geminiteam2024gemini] LLM-as-a-judge 0 100%
+RvLLM 100% 100%
+SelfCheckGPT 100% 0%
+DeepSeek-V3 LLM-as-a-judge 0 100%
+RvLLM 100% 93.2%
+Table 11: Performance of different methods for inequality solving tasks in case study 3.
+Target LLMs Methods TPR TNR Token Usage Per Task
+SelfCheckGPT 100% 0 9850
+Qwen-turbo [yang2024qwen2] LLM-as-a-judge 100% 0 1466
+RvLLM 18.2% 94.4% 4805
+SelfCheckGPT 100% 0 10373
+GPT 4.1 mini [achiam2023gpt] LLM-as-a-judge 100% 0 2479
+RvLLM 50.0% 84.2% 5683
+Appendix F Experiment statistical significance
+The only source of variability in RvLLM arises from the non-deterministic outputs of the perception agent.
+To reduce this variability while maintaining the effectiveness of its text-understanding capability during the interpretation abstraction phase, we set the temperature to 0.5 and conduct preliminary runs until the outputs stabilize.
+This process ensures that the reported experimental results exhibit negligible variance.
+Appendix G Broader impacts
+Integrating domain-specific knowledge into LLM verification, as proposed in this work, enhances the reliability of LLMs in high-stakes settings.
+Through a general specification language ESL and a runtime verification framework RvLLM, domain experts can encode structured knowledge to systematically detect and mitigate inconsistencies in model outputs.
+This improves response accuracy in specialized domains—such as legal compliance or mathematical reasoning—and supports the deployment of LLMs in safety-critical systems that require stronger guarantees.
+Nonetheless, risks remain.
+The effectiveness of RvLLM relies on both the perception agent and the quality of human-crafted specifications; deficiencies in either may introduce bias or error.
+Thus, while our framework provides a promising step toward more robust LLM behavior, it should be viewed as a complementary safeguard rather than a complete solution to the broader challenges of ensuring reliable LLM outputs.